### v2

# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

Abstract
The Poisson compound decision problem is a long-standing problem in statistics, where empirical Bayes methodologies are commonly used to estimate Poisson’s means in static or batch domains. In this paper, we study the Poisson compound decision problem in a streaming or online domain. Adopting a quasi-Bayesian approach, referred to as Newton’s algorithm, we obtain a sequential estimate that is easy to evaluate, computationally efficient, and maintain a constant per-observation computational cost as data accumulate. Asymptotic frequentist guarantees of this estimate are established, showing consistency and asymptotic optimality, where the latter is understood as vanishing excess Bayes risk or regret. We demonstrate the effectiveness of our methodology through empirical analysis on synthetic and real data, with comparisons to existing approaches.
Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data
1 Introduction
Given $n\geq 1$ observations modeled as independent Poisson random variables $Y_{1},\ldots,Y_{n}$ with means $\theta_{1},\ldots,\theta_{n}$, respectively, the Poisson compound decision problem refers to the estimation of $(\theta_{1},\ldots,\theta_{n})$ under the squared-error loss. Empirical Bayes is a general approach to handle compound decision problems (Robbins, 1951, 1956; Zhang, 2003). Assuming the $\theta_{i}$’s to be i.i.d. as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ is the posterior mean or Bayes estimate, namely for $y\in\mathbb{N}_{0}$
$\hat{\theta}_{G}(y)=E_{G}[\theta_{i}\mid Y_{i}=y]=\frac{\int_{\Theta}\theta\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}{\int_{\Theta}\frac{\text{e}^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)}=(y+1)\frac{p_{G}(y+1)}{p_{G}(y)},$ (1)
where $p_{G}$ is the probability mass function of the $Y_{i}$’s. The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either the marginal $p_{G}$, a strategy known as $f$-modeling, or the prior ${G}$, a strategy known as $g$-modeling (Efron, 2014).
1.1 Background and motivation
A prominent example of $f$-modeling is Robbins’ method (Robbins, 1956), which estimates $p_{G}$ in (1) with the empirical distribution of the $Y_{i}$’s (Brown et al., 2013; Polyanskiy and Wu, 2021; Jana et al., 2023); see Efron (2014, Section 3 and Section 4). Although widely used in practice for its computational simplicity, $f$-modeling lacks robustness and numerical stability; in fact, it is easily derailed by “outliers” in the data, namely data points that appear only a few times, causing estimators to take exceptionally small or large values (Efron and Hastie, 2021; Shen and Wu, 2024).
Examples of $g$-modeling include estimates of $G$ in (1) based on maximum likelihood and minimum distance methods (Jiang and Zhang, 2009; Jana et al., 2024); see Efron (2014, Section 5). A fully Bayesian approach is also possible by placing a hyperprior on $G$, a strategy known as Bayes empirical Bayes (Deely and Lindley, 1981; Efron, 2019). While $g$-modeling yields more accurate estimates than $f$-modeling, as well as better incorporating prior knowledge, it is computationally demanding, particularly in nonparametric and high-dimensional settings (Jana et al., 2023).
The Poisson compound decision problem offers a flexible framework for modeling count data, with successful applications in settings where static or batch processing has traditionally sufficed, such as microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024). However, the rise of streaming data in some of these fields, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, calls for extending the Poisson compound decision problem to the online setting, where timely analysis is critical for responsive and informed decision-making. Although the existing empirical Bayes methods can, in principle, be adapted to streaming data, they often suffer from computational inefficiencies, particularly when involving optimization or posterior computations.
1.2 Preview of our contributions
We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem. Starting from an initial guess $G_{0}$ for $G$, for $n\geq 1$ our approach updates $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008). This recursive procedure is regarded as a quasi-Bayesian learning model, yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model (Fortini and Petrone, 2020). A quasi-Bayes empirical Bayes estimate of $\hat{\theta}_{G}(y)$ is then obtained from (1) by replacing $G$ with $G_{n}$. Besides being easy to evaluate, the estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient, and maintain a constant per-observation computational cost as data accumulate. Considering Newton’s algorithm as a learning model allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, for which we build an asymptotic credible interval via the central limit theorem.
Asymptotic frequentist guarantees are established for $\hat{\theta}_{G_{n}}(y)$, assuming that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$. We provide sufficient conditions on $G^{\ast}$ to ensure consistency and asymptotic optimality of $\hat{\theta}_{G_{n}}(y)$ as $n\rightarrow+\infty$ , where the latter is understood as vanishing excess Bayes risk or regret. Under additional assumptions on Newton’s algorithm, the rate at which the regret converges to zero is also derived.
We validate our methodology empirically on synthetic and real data. For fixed $n\geq 1$, the estimate $\hat{\theta}_{G_{n}}(y)$ consistently outperforms Robbins’ method and performs competitively with the nonparametric maximum likelihood method and the minimum squared Hellinger distance method. It also performs on par with a parametric Bayes empirical Bayes method with a Gamma hyperprior.
2 Quasi-Bayes empirical Bayes
For a stream of observations $(Y_{n})_{n\geq 1}$, we propose a sequential $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies the distribution of $(Y_{n})_{n\geq 1}$ through the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$. As in the Bayesian framework, the model is conceived here as a learning model and not as a data-generating process (Fortini and Petrone, 2025).
Letting $(\mathcal{G}_{n})_{n\geq 1}$ denote the natural filtration of $(Y_{n})_{n\geq 1}$, for each $n\geq 0$ the conditional distribution of $Y_{n+1}$ given $\mathcal{G}_{n}$ is modeled according to a mixture of the Poisson kernel $k(\cdot\mid\theta)$, with $\theta$ ranging over a sufficiently fine and wide grid $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ for $j=1,\dots,d-1$, i.e.,
$p_{g_{n}}(y)=\mathds{P}(Y_{n+1}=y\mid\mathcal{G}_{n})=\sum_{j=1}^{d}k(y\mid\vartheta_{j})g_{n}(\vartheta_{j})\quad y\in\mathbb{N}_{0},$ (2)
where the mixing distribution $g_{n}$ is defined recursively by means of the following Newton’s algorithm as
$g_{n+1}(\vartheta_{j})=(1-\alpha_{n+1})g_{n}(\vartheta_{j})+\alpha_{n+1}\frac{k(Y_{n+1}\mid\vartheta_{j})g_{n}(\vartheta_{j})}{\sum_{i=1}^{d}k(Y_{n+1}\mid\vartheta_{i})g_{n}(\vartheta_{i})}\qquad j=1,\ldots,d,$ (3)
given an initial guess $g_{0}$ and the $\alpha_{n}$’s in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$. From (3), the learning model updates $g_{n}$ as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, between $g_{n}$ and its posterior distribution based on $Y_{n+1}$. The sequence $(\alpha_{n})_{n\geq 1}$ is referred to as the learning rate, for which a standard choice is $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$.
The following proposition outlines an effective procedure for constructing the grid $\Theta_{d}$ used in (2)-(3). Beyond serving as a practical tool for implementing the sequential $g$-modeling methodology, this construction also plays a central role in establishing asymptotic frequentist guarantees. We denote by $\text{KL}(p||q)$ the Kullback-Leibler divergence between the probability distributions $p$ and $q$.
Proposition 2.1.
Let the prior $G$ in (1) be such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ define the equally spaced grid $\Theta_{\eta}=\{\vartheta_{i}=i\eta\text{ : }i=1,\ldots,d_{\eta}\}$ with $d_{\eta}=\inf\{n\in\mathbb{N}\text{ : }n>\eta^{-1}\text{ and }n^{1-k}\log(n\eta)m_{k}\leq\eta^{k}\}$. If $\mathcal{S}_{\eta}$ is the set of all the probability mass functions on $\Theta_{\eta}$, then $g_{\eta}=\text{argmin}_{f\in\mathcal{S}_{\eta}}\text{KL}(p_{G}||p_{f})$ is unique and $\text{KL}(p_{G}||p_{g_{\eta}})<2\eta$.
See C.1 for the proof of Proposition 2.1. The probability mass function $g_{\eta}$ in Proposition 2.1 is referred to as the $d_{\eta}$-KL-discretization of a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$. This form of discretization is of independent interest in $g$-modeling approaches to the Poisson compound decision problem (Jana et al., 2024).
2.1 Quasi-Bayes properties
We show that learning model defined by (2)-(3) is quasi-Bayesian in the sense of Fortini and Petrone (2020), yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model.
Theorem 2.2.
[…]
See C.2 for the proof of Theorem 2.2. Under the learning model (2)-(3), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are considered conditionally identically distributed given $\mathcal{G}_{n}$. From Berti et al. (2004, Theorem 2.2), this implies that the $Y_{n}$’s are asymptotically i.i.d., given the random limit $\tilde{g}$, with law $p_{\tilde{g}}$, which fits within the quasi-Bayesian framework of Fortini and Petrone (2020). In this perspective, $p_{\tilde{g}}$ represents the random distribution that would be assigned if we had access to an infinitely large sample, which makes $\tilde{g}$ the parameter of interest. Accordingly, (1) becomes
$\hat{\theta}_{\tilde{g}}(y)=(y+1)\frac{p_{\tilde{g}}(y+1)}{p_{\tilde{g}}(y)}\qquad y\in\mathbb{N}_{0}.$ (4)
Differently from the Bayesian framework, which posits a prior distribution on $g$ such that the $Y_{n}$’s are i.i.d. given $g$, the quasi-Bayesian approach avoids the computational cost of recomputing the posterior distribution with each new observation, thereby offering substantial computational efficiency.
2.2 Sequential estimation and credible intervals
Under the learning model (2)-(3), an estimate of $\hat{\theta}_{\tilde{g}}(y)$ is obtained by replacing $\tilde{g}$ in (4) with $g_{n}$, i.e.
$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0},$ (5)
which is referred to as quasi-Bayes empirical Bayes. As a corollary of Theorem 2.2, for every $y\in\mathbb{N}_{0}$ it holds that $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s. to $\hat{\theta}_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. The following central limit theorem allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, building an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$.
Theorem 2.3.
Let (2)-(3) hold with $(\alpha_{n})_{n\geq 1}$ non increasing and $\sum_{n\geq 1}(\alpha_{n}^{2}/\sum_{k\geq n}\alpha_{n}^{2})^{2}<+\infty$, and let $b_{n}=(\sum_{k\geq n}\alpha_{k}^{2})^{-1}$. For $y\in\mathbb{N}_{0}$, the conditional distribution of $b_{n}^{1/2}(\hat{\theta}_{\tilde{g}}(y)-\hat{\theta}_{g_{n}}(y)$, given $\mathcal{G}_{n}$, converges weakly $\mathds{P}$-a.s. to a centered (i.e., zero-mean) Gaussian kernel with the random variance
$W_{\tilde{g}}(y)=\hat{\theta}^{2}_{\tilde{g}}(y)E_{Z\sim p_{\tilde{g}}}\left[\left(\sum_{j=1}^{d}\frac{k(Z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(Z)}\left(\frac{k(y+1\mid\vartheta_{j})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{j})}{p_{\tilde{g}}(y)}\right)\right)^{2}\right].$ (6)
See C.3 for the proof of Theorem 2.3. If $W_{g_{n}}(y)$ is defined as in (6), with $g_{n}$ in the place of $\tilde{g}$, then for every $y\in\mathbb{N}_{0}$ it holds that $W_{g_{n}}(y)$ converges $\mathds{P}$-a.s. to $W_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately a Gaussian distribution with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$. This result provides an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
A sequential $g$-modeling methodology may be also considered from a frequentist perspective, where $(Y_{n})_{n\geq 1}$ are i.i.d. from a Poisson mixture model with a “true” mixing distribution $g^{\ast}$, and Newton’s algorithm is employed as a consistent estimate of $g^{\ast}$ (Tokdar et al., 2009; Martin and Tokdar, 2009). While this perspective leads to the same estimate $\hat{\theta}_{g_{n}}(y)$ as (5), constructing asymptotic confidence intervals remains out of reach, as the stochastic approximation framework underlying Newton’s method (Martin and Ghosh, 2008) does not apply to Gaussian central limit theorems.
2.3 Frequentist guarantees
The “true” Poisson compound model assumes that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$ on $\Theta\subseteq\mathbb{R}^{+}$. If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ in (1) with the oracle $\hat{\theta}_{G^{*}}(y)$. Instead, not knowing the oracle prior $G^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (5). Here, Proposition 2.1 is leveraged in order to establish asymptotic frequentist guarantees of $\hat{\theta}_{g_{n}}(y)$ under the “true” model for $(Y_{n})_{n\geq 1}$, demonstrating consistency and asymptotic optimality in terms of vanishing excess Bayes risk or regret.
Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ that satisfies the assumptions of Proposition 2.1, and, for a fixed $\eta>0$, let $g^{\ast}_{\eta}$ be its $d_{\eta}$-KL-discretization. We define the regret of using $\hat{\theta}_{g_{n}}(y)$ instead of $\hat{\theta}_{g_{\eta}^{*}}(y)$ as
$\text{Regret}(g_{n},g_{\eta}^{*})=E_{G^{*}}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g_{\eta}^{*}}(Y_{n}))^{2}=\int_{\Theta}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}k(y\mid\theta)G^{\ast}(\mathrm{d}\theta)$ (7)
with $E_{G^{*}}$ being the expected value with respect to the “true” model; see Efron (2019, Section 3) for details on (7). We denote by $\mathds{P}_{G^{*}}$ the probability measure under which the “true” model for $(Y_{n})_{n\geq 1}$ holds true.
Theorem 2.4.
Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G^{\ast}}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ let $g^{\ast}_{\eta}$ be the $d_{\eta}$-KL-discretization of $G^{\ast}$ on $\Theta_{\eta}$. For each $n\geq 1$, let $g_{n}(\vartheta_{j})$ be as in (3) for $\vartheta_{j}\in\Theta_{\eta}$, with $g_{0}(\vartheta_{j})>0$ for every $j=1,\dots,d_{\eta}$, and with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$. If $\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{\ast}}(y)<+\infty$ then, for every $y\in\mathbb{N}_{0}$,
$\hat{\theta}_{g_{n}}(y)\rightarrow\hat{\theta}_{g^{\ast}_{\eta}}(y)\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (8)
Furthermore, if $\gamma<1$ then for every $\delta<2-1/\gamma$
$\text{Regret}(g_{n},g_{\eta}^{\ast})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (9)
See C.4 for the proof of Theorem 2.4. Under the “true” model for the observations $Y_{n}$’s, with the oracle prior $G^{\ast}$ belonging to the class of light-tail distributions, Theorem 2.4 establishes that $\hat{\theta}_{g_{n}}(y)$ is a consistent (8) and asymptotically optimal (9) estimate of $\hat{\theta}_{g^{\ast}_{\eta}}(y)$. Although Theorem 2.4 is stated with respect to the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$, alternative learning rates may also be considered.
3 Numerical illustrations
We validate our methodology empirically, by comparing the quasi-Bayes empirical Bayes (QB-EB) estimate $\hat{\theta}_{g_{n}}$, at fixed $n\geq 1$, with estimates from: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981). NP-EB uniquely yields a closed-form estimate by replacing $p_{G}$ in (1) with the empirical distribution, while NP-ML, NP-MD and P-EB require exact or approximate optimization algorithms. NP-ML and NP-MD estimates follow the vertex direction algorithm Jana et al. (2024, Algorithm 1), which, as QB-EB, relies on a discretization of $G$ over a finite grid. P-EB estimates use Deely and Lindley (1981, Equation 4.4), with hyperparameters of $G$ estimated by maximum (marginal) likelihood. Performance is assessed via root mean squared error (RMSE) and mean absolute deviation (MAD).
3.1 Synthetic data
[…]
Figure 1: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and QB-EB (blue).
Table 1: Values of RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
[…]
3.2 Real data
We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/. It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts. For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting: fast retweeting is often indicative of automated or bot-driven activity, as opposed to human behavior. The underlying retweet intensity can thus be viewed as a latent measure of a tweet’s instantaneous virality, or more generally its propensity to trigger immediate attention. Identifying tweets with unusually high intensity in real time could help detect automated activity or early-stage virality. Since these counts arrive rapidly in a streaming fashion, they motivate the need for a sequential, online estimation method such as the proposed QB-EB.
Table 2 summarizes the empirical distribution of fast retweet counts, e.g., 40,259 tweets received none, 28,339 received one, etc. We apply NP-EB, NP-ML, NP-MD, P-FB, and QB-EB to estimate the intensity for each tweet, quantifying its fast-retweet potential. Results are shown in Table 2.
Table 2: Counts $n_{y}$ of the number of tweets with $y$ fast retweets, and corresponding estimates.
$y$ 0 1 2 3 4 5 6 7
$n_{y}$ 40,259 28,339 21,581 16,479 12,130 9,238 7,193 5,464
NP-EB 0.704 1.523 2.291 2.944 3.808 4.672 5.317 6.324
NP-ML 0.391 2.134 2.931 3.697 4.459 5.329 6.334 7.424
NP-MD 0.750 2.198 3.142 4.149 4.986 5.694 6.390 7.163
P-EB 0.445 1.335 2.226 3.116 4.007 4.897 5.788 6.678
QB-EB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
We refer to D.2 for another application of the Twitter dataset, with respect to the followers generated by retweets, and for an application to the benchmark accident dataset (Efron and Hastie, 2021, Table 6.1).
4 Discussion
An extension of the sequential $g$-modeling methodology to the multidimensional Poisson compound decision problem (Brown et al., 1985) is presented in Section E. While our focus is on the Poisson setting, the methodology is readily extendable to other discrete linear exponential families (Banerjee et al., 2021), as well as to the Gaussian setting (Jiang and Zhang, 2009; Soloff et al., 2024).
Appendix
B Preliminaries
B.1 Notation
[…]
B.2 Almost sure conditional convergence.
Given a probability space $(\Omega,\mathcal{F},\mathds{P})$ and a Polish space $\mathbb{X}$ with its Borel sigma-algebra $\mathcal{X}$, a kernel on $\mathbb{X}$ is a function $K:\mathcal{X}\times\Omega$ satisfying: i) for every $\omega\in\Omega$, $K(\cdot,\omega)$ is a probability measure on $\mathcal{X}$; ii) for each $B\in\mathcal{X}$, the function $K(B,\cdot)$ is measurable with respect to $\mathcal{F}$. A kernel $K$ on $\mathbb{R}$ is called a Gaussian kernel if $K(\cdot,\omega)$ is a Gaussian distribution with mean $\mu(\omega)$ and variance $\sigma^{2}(\omega)$, where $\mu$ and $\sigma^{2}$ are random variables defined on $(\Omega,\mathcal{F},\mathds{P})$. We denote the Gaussian kernel by $\mathcal{N}(\mu,\sigma^{2})$ and interpret the Gaussian distribution with zero variance as the degenerate law centered on the mean. A kernel $K$ on the Polish space $C(\mathbb{R})$, with the topology of uniform convergence on compact sets, is called a Gaussian process kernel if $K(\cdot,\omega)$ is the distribution of a Gaussian process in $C(\mathbb{R})$, with mean function $\mu(\omega)\in C(\mathbb{R})$ and covariance function $R(\omega)\in C(\mathbb{R}\times\mathbb{R})$. If $\mu(\omega)=0$ for all $\omega$, then the Gaussian process kernel is said to be centered. We extend the definition of almost sure conditional convergence for real sequences (Crimaldi, 2009), to sequences taking values in Polish spaces.
Definition B.1.
Let $(X_{n})$ be a sequence of random variables adapted to a filtration $(\mathcal{G}_{n})$ and taking values in a Polish space $\mathbb{X}$ with its Borel sigma-algebra $\mathcal{X}$. For every $n\geq 0$, let $K_{n}$ denote a regular version of the conditional distribution of $X_{n+1}$, given $\mathcal{G}_{n}$. If there exists a kernel $K$ such that the sequence $(K_{n}(\cdot,\omega))_{n}$ converges weakly to $K(\cdot,\omega)$ for almost every $\omega\in\Omega$, then we say that the sequence $(X_{n})$ converges to $K$ in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{G}_{n})$.
The results are the main tool in the proof of Theorem 2.3.
Theorem B.2 (Crimaldi (2009), Theorem A.1).
For each $n\geq 1$, let $(M_{n,j})_{j\geq 1}$ be a real valued martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 1}$, satisfying $M_{n,0}=0$, and converging in $L^{1}$ to a random variable $M_{n,\infty}$. Set
$X_{n,j}=M_{n,j}-M_{n,j-1}\mbox{ for }j\geq 1,\quad U_{n}=\sum_{j\geq 1}X_{n,j}^{2},\quad X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|.$
Assume that
  • (a)
    $(X_{n}^{*})_{n}$ is dominated in $L^{1}$ and converges to zero a.s.
  • (b)
    $(U_{n})_{n}$ converges a.s. to a non-negative random variable $U$.
Then the sequence $(M_{n,\infty})$ converges to the Gaussian kernel $\mathcal{N}(0,U)$ in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{F}_{n,1})$.
Remark B.3.
Requesting that the almost sure limit of $(U_{n})$ is non-negative guarantees the well-definedness of the kernel $\mathcal{N}(0,U)$ for every $\omega\in\Omega$. Theorem A.1 in Crimaldi (2009) specifies that $U$ should be a “positive random variable”. Although not explicitly stated, this requirement should be understood as $U\geq 0$, as becomes evident upon a careful examination of the proof.
Theorem B.4 (Crimaldi et al. (2016), Lemma 4.1).
Let $(Z_{n})$ be a sequence of real valued random variables adapted to the filtration $(\mathcal{G}_{n})$ and such that $E(Z_{n+1}\mid\mathcal{G}_{n})\rightarrow Z$ a.s. for some random variable $Z$. Moreover, let $(a_{n})$ and $(b_{n})$ be sequences of real numbers such that
$b_{n}\uparrow+\infty,\quad\sum_{k=1}^{\infty}\frac{E(Z_{k}^{2})}{a_{k}^{2}b_{k}^{2}}<+\infty.$
Then we have:
  • (a)
    If $\frac{1}{b_{n}}\sum_{k=1}^{n}\frac{1}{a_{k}}\rightarrow c$ for some constant $c$, then $\frac{1}{b_{n}}\sum_{k=1}^{n}\frac{Z_{k}}{a_{k}}\rightarrow cZ$ a.s.
  • (b)
    If $b_{n}\sum_{k\geq n}\frac{1}{a_{k}b_{k}^{2}}\rightarrow c$ for some constant $c$, then ${b_{n}}\sum_{k\geq n}\frac{Z_{k}}{a_{k}b_{k}^{2}}\rightarrow cZ$ a.s.
B.3 Delta method for almost-sure conditional convergence
In this section, we extend the delta method to almost sure conditional convergence. Let $\mathbf{h}$ be a function defined on a convex set $\Delta\in\mathcal{B}(\mathbb{R}^{d})$ with values in $\mathbb{R}^{p}$. The function $\mathbf{h}$ is said to be uniformly differentiable on $\Delta$ if there exists, for every $\mathbf{g}^{*}\in\Delta$ a linear function $\mathbf{D}\mathbf{h}_{\mathbf{g}^{*}}:\Delta\rightarrow\mathbb{R}^{p}$ such that the following condition holds: for every $\epsilon>0$ there exists $\delta$ such that
$\frac{||\mathbf{h}(\mathbf{g})-\mathbf{h}(\mathbf{g}^{*})-\mathbf{D}\mathbf{h}_{\mathbf{g}^{*}}(\mathbf{g}-\mathbf{g}^{*})||}{||\mathbf{g}-\mathbf{g}^{*}||}<\epsilon\quad\mbox{ for every }\quad\mathbf{g}\neq\mathbf{g}^{*},||\mathbf{g}-\mathbf{g}^{*}||<\delta.$ (B.1)
Theorem B.5.
Let $(\mathbf{g}_{n})$ be a sequence of random vectors taking values in a convex set $\Delta\in\mathcal{B}(\mathbb{R}^{d})$ with values in $\mathbb{R}^{p}$, adapted to the filtration $(\mathcal{G}_{n})$, and let $\tilde{\mathbf{g}}$ be a random vector taking values in $\Delta$. Assume that there exist an increasing and diverging sequence $(r_{n})$ of positive numbers and a kernel $K(\cdot,\omega)$ tight with respect to $\omega$, $P$-a.s., such that $r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to $K(\cdot,\omega)$. If $\mathbf{h}$ is a uniformly differentiable function on $\Delta$, then $r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence to the kernel $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\cdot),\omega)$.
The proof is split into several lemmas.
Lemma B.6.
Under the assumptions of Theorem B.5, for every $\epsilon>0$
$P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)\rightarrow 0\quad P-a.s.$
as $n\rightarrow\infty$, where $P_{n}(\cdot)=P(\cdot\mid\mathcal{G}_{n})$.
Proof.
Fix $\epsilon>0$ and let $\delta$ be such that (B.1) holds. Then
$P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-D\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)$ $\leq P_{n}(||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq\delta)\leq P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq r_{n}\delta).$
Fix $\eta$, and let $M$ be such that $K([-M,M]^{d},\omega)>1-\eta$ for almost every $\omega$. Then, there exists $n_{0}=n_{0}(\omega)$ such that, for $\omega$ in a set of probability one, and $n>n_{0}(\omega)$,
$P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)<2\eta.$
It follows that
$\limsup_{n}P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)<2\eta\quad\quad P-a.s.$
The thesis follows from the arbitrariness of $\eta$. ∎
Lemma B.7.
Under the assumptions of Theorem B.5, for every $\epsilon>0$,
$P_{n}\left(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon\right)\rightarrow 0\quad P-a.s.$
as $n\rightarrow\infty$.
Proof.
It can be proved as in Lemma B.6 that, for every $\eta>0$, there exists $n_{0}=n_{0}(\omega)$ such that, for $\omega$ in a set of probability one and $n\geq n_{0}(\omega)$,
$P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)<\eta.$
Thus, for $n>n_{0}(\omega)$, it holds, $P$-a.s. that
$P_{n}\left(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon\right)(\omega)$
$\quad\leq P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}\;r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||>\epsilon\right)(\omega)$
$\quad\leq P_{n}(r_{n}||\mathbf{g}_{n}-\tilde{\mathbf{g}}||\geq M)(\omega)+P_{n}\left(M\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\epsilon\right)(\omega)$
$\quad\leq\eta+P_{n}\left(\frac{||\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}})-\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||}{||\mathbf{g}_{n}-\tilde{\mathbf{g}}||}>\frac{\epsilon}{M}\right)(\omega).$
The thesis follows from Lemma B.6 and the arbitrariness of $\eta$.
of Theorem B.5.
Let $N^{c}$ be an event such that $P(N^{c})=1$, $\{K(\cdot,\omega):\omega\in N^{c}\}$ is tight, $P_{n}(r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}})\in\cdot)(\omega)\rightarrow K(\cdot,\omega)$ and $P_{n}(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{g}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon)(\omega)\rightarrow 0$ for every $\epsilon>0$. Fix $\omega\in N^{c}$ and let $A\in\mathcal{B}(\mathbb{R}^{p})$ be such that $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\partial A),\omega)=0$. For every $\epsilon>0$, let $A^{\epsilon}=\{\mathbf{x}\in\mathbb{R}^{p}:{\rm d}(\mathbf{x},A)\leq\epsilon\}$. Then
$\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)$
$\leq\limsup_{n}P_{n}(||r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))-r_{n}\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(\mathbf{g}_{n}-\tilde{\mathbf{g}})||>\epsilon)(\omega)$
$+\limsup_{n}P_{n}(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}(r_{n}(\mathbf{g}_{n}-\tilde{\mathbf{g}}))\in A^{\epsilon})(\omega)$
$\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega).$
Analogously,
$\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A^{c})(\omega)\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}((A^{c})^{\epsilon}),\omega).$
Thus,
$\liminf_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)\geq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega),$
where $A_{\epsilon}$ is the complement of $(A^{c})^{\epsilon}$. Thus, for every $\epsilon>0$,
$K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega)$ $\leq\liminf_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)$
$\leq\limsup_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)$
$\leq K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega).$
Since $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(\partial A,\omega)=0$, then $K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A^{\epsilon}),\omega)-K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A_{\epsilon}),\omega)\rightarrow 0$, as $\epsilon\rightarrow 0$. Thus $\lim_{n}P_{n}(r_{n}(\mathbf{h}(\mathbf{g}_{n})-\mathbf{h}(\tilde{\mathbf{g}}))\in A)(\omega)=K(\mathbf{D}\mathbf{h}_{\tilde{\mathbf{g}}}^{-1}(A),\omega).$
C Proofs for Section 2
C.1 Proof of Proposition 2.1
We first prove that there exists a probability mass function $f\in\mathcal{S}_{\eta}$ such that $KL(p_{G}||p_{f})<2\eta.$ Since the $k$-moment of the Poisson distribution with parameter $\lambda$ is greater than $\lambda^{k}$, then the assumption $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ implies that
$\int_{0}^{\infty}\theta^{k}G(\mathrm{d}\theta)\leq m_{k}.$
For every $i=1,\dots,d_{\eta}$, let $w_{i}=G((\vartheta_{i-1},\vartheta_{i}])$ with $\vartheta_{0}=0$. Let $F$ be the probability measure
$F=\sum_{i=1}^{d_{\eta}}w_{i}\delta_{\vartheta_{i}}+G(\vartheta_{d_{\eta}},+\infty)\delta_{\vartheta_{d_{\eta}}},$
and let $f$ denote its density with respect to the counting measure on $\Theta_{\eta}$. Then
$F=\sum_{i=1}^{d_{\eta}}\delta_{\vartheta_{i}}\int_{I_{i}}G(\mathrm{d}\theta)+\delta_{\vartheta_{d_{\eta}}}\int_{I_{d_{\eta}+1}}G(\mathrm{d}\theta),$
where $I_{i}=(\vartheta_{i-1},\vartheta_{i}]$ for $i=1,\dots,d_{\eta}$ and $I_{d_{\eta}+1}=(\vartheta_{d_{\eta}},+\infty)$. Since the Kullback-Leibler divergence is convex, then, denoting $M=d_{\eta}\eta$, we can write that
$KL(p_{G}||p_{f})$ $\leq\sum_{y\in\mathbb{N}_{0}}\sum_{i=1}^{d_{\eta}}\int_{I_{i}}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-\vartheta_{i}}\vartheta_{i}^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)+\sum_{y\in\mathbb{N}_{0}}\int_{M}^{\infty}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-M}M^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
$<\int_{0}^{M}\eta G(\mathrm{d}\theta)+\int_{M}^{\infty}\sum_{y\in\mathbb{N}_{0}}y\log(\theta)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
$\leq\eta+\int_{M}^{\infty}\theta\log(\theta)G(\mathrm{d}\theta)$
$\leq\eta+\frac{\log M}{M^{k-1}}\int_{M}^{\infty}\theta^{k}G(\mathrm{d}\theta)$
$\leq 2\eta.$
We now prove that the minimization problem $\mbox{\sl argmin}_{f\in\mathcal{S}_{\eta}}KL(p_{G}||p_{f})$ has a unique solution. Representing each probability mass function $f$ with support $\Theta_{\eta}$ as a vector $\mathbf{f}$ taking values in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a strictly convex function defined on a convex and compact set. Thus the global minimum $\mathbf{g}_{\eta}$ exists and is unique. $\square$
C.2 Proof of Theorem 2.2
For every $n\geq 0$ and every $j=1,\dots,d$,
$E(g_{n+1}(\vartheta_{j})\mid\mathcal{G}_{n})=(1-\alpha_{n+1})g_{n}(\vartheta_{j})+\alpha_{n+1}\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{j})g_{n}(\vartheta_{j})}{\sum_{i=1}^{d}k(y\mid\vartheta_{i})g_{n}(\vartheta_{i})}p_{g_{n}}(y)=g_{n}(\vartheta_{j}).$
Thus, the sequence $(g_{n}(\vartheta_{j}))$ is a bounded martingale with respect to the filtration $(\mathcal{G}_{n})$. Denoting by $\tilde{g}(\vartheta_{j})$ its $\mathds{P}$-a.s. limit, we have that $\tilde{g}$ is a probability mass function on $\Theta_{d}$, $g_{n}(\vartheta_{j})=E(\tilde{g}(\vartheta_{j})\mid\mathcal{G}_{n})$ for every $n\geq 0$ and $j=1,\dots,d$, and, for every $y\in\mathbb{N}_{0}$, $p_{g_{n}}(y)$ converges $\mathds{P}$-a.s. to $p_{\tilde{g}}(y)$. Since $\tilde{g}$ is a probability mass function on $\Theta_{d}$, then $p_{\tilde{g}}$ is a probability mass function on $\mathbb{N}_{0}$. Moreover, for every $y\in\mathbb{N}_{0}$,
$E(p_{g_{n+1}}(y)\mid\mathcal{G}_{n})=E(\sum_{j=1}^{d}k(y\mid\vartheta_{j})g_{n+1}(\vartheta_{j})\mid\mathcal{G}_{n})=p_{g_{n}}(y).$
Thus, $(p_{g_{n}}(y))$ is also a martingale with respect to $\mathcal{G}_{n}$, and $p_{g_{n}}(y)=E(p_{\tilde{g}}(y)\mid\mathcal{G}_{n})$ for every $y\in\mathbb{N}_{0}$ and $n\geq 0$. It follows that, for every $y\in\mathbb{N}_{0}$, every $n\geq 0$ and every $k\geq 1$,
$\mathds{P}(Y_{n+k}=y\mid\mathcal{G}_{n})=E(\mathds{P}(Y_{n+k}=y\mid\mathcal{G}_{n+k-1}\mid\mathcal{G}_{n})=E(p_{g_{n+k-1}}(y)\mid\mathcal{G}_{n})=p_{g_{n}}(y)=E(p_{\tilde{g}}(y)\mid\mathcal{G}_{n}),$
[…]
C.3 Proof of Theorem 2.3
For every $z\in\mathbb{N}_{0}$ let $\tilde{g}(\vartheta_{j}\mid z)=\frac{k(z\mid\vartheta_{j})\tilde{g}(\vartheta_{j})}{p_{\tilde{g}}(z)}$. Let us define the random matrices $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ ($n\geq 0$) with $(i,j)$ element
$\mathbf{V}_{\tilde{g},i,j}=\sum_{z\in\mathbb{N}_{0}}\tilde{g}(\vartheta_{i}\mid z)\tilde{g}(\vartheta_{j}\mid z)p_{\tilde{g}}(z)-\tilde{g}(\vartheta_{i})\tilde{g}(\vartheta_{j}).$ (C.2)
$\mathbf{V}_{n,i,j}=\sum_{z\in\mathbb{N}_{0}}g_{n}(\vartheta_{i}\mid z)g_{n}(\vartheta_{j}\mid z)p_{g_{n}}(z)-g_{n}(\vartheta_{i})g_{n}(\vartheta_{j}),$ (C.3)
for $i,j=1,\dots,d-1$.
Lemma C.1.
Let $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ be defined as in (C.2) and (C.3), respectively. Then, $\mathds{P}$-a.s., $(\mathbf{V}_{n})$ is a sequence of uniformly bounded, positive definite random matrices, converging to $\mathbf{V}_{\tilde{g}}$ as $n\rightarrow\infty$.
Proof.
Since, for every $i,j$
$|\mathbf{V}_{n,i,j}|\leq 1,$
then the sequence $(\mathbf{V}_{n})$ is uniformly bounded. Moreover for every vector $\mathbf{u}\in\mathbb{R}^{d-1}$,
$\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}$ $=\sum_{i,j=1}^{d-1}\sum_{z\in\mathbb{N}_{0}}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))\mathbf{u}_{j}(g_{n}(\vartheta_{j}\mid z)-g(\vartheta_{j}))p_{g_{n}}(z)$
$=\sum_{z\in\mathbb{N}_{0}}\left(\sum_{i=1}^{d-1}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))\right)^{2}p_{g_{n}}(z)\geq 0,$
$\mathds{P}$-a.s. Hence, $\mathds{P}$-a.s., $\mathbf{V}_{n}$ is positive semi-definite. We now prove that it is $\mathds{P}$-a.s. positive definite. By contradiction, suppose that there exists a random vector $\mathbf{u}$ such that for $\omega$ in a set $N^{c}$ with positive probability $\mathbf{u}(\omega)\neq\mathbf{0}$ and $\mathbf{u}^{T}(\omega)\mathbf{V}_{n}(\omega)\mathbf{u}(\omega)=0$. Fix $\omega\in N^{c}$ and, without loss of generality, assume that $p_{g_{n}(\omega)}(z)>0$ for every $z\in\mathbb{N}_{0}$. Then, we can write (omitting $\omega$ for simplicity)
$\sum_{i=1}^{d-1}\mathbf{u}_{i}(g_{n}(\vartheta_{i}\mid z)-g_{n}(\vartheta_{i}))=0,$
for every $z\in\mathbb{N}_{0}$. It follows that for every $z\in\mathbb{N}_{0}$,
$\sum_{i=1}^{d-1}\mathbf{u}_{i}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})\sum_{i=1}^{d-1}\mathbf{u}_{i}g_{n}(\vartheta_{i})$
Define $u(\vartheta_{i})=\mathbf{u}_{i}$ for $i=1,\dots,d-1$ and $u(\vartheta_{d})=0$. Then
$\sum_{i=1}^{d}u(\vartheta_{i})e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{-\vartheta_{i}}\vartheta_{i}^{z}g_{n}(\vartheta_{i})\sum_{i=1}^{d}u(\vartheta_{i})g_{n}(\vartheta_{i}),$
which implies that, for every $k=0,1,\dots$,
$\sum_{i=1}^{d}u(\vartheta_{i})e^{k\vartheta_{i}}g_{n}(\vartheta_{i})=\sum_{i=1}^{d}e^{k\vartheta_{i}}g_{n}(\vartheta_{i})\sum_{i=1}^{d}u(\vartheta_{i})g_{n}(\vartheta_{i}).$ (C.4)
To demonstrate that this leads to a contradiction, consider a random variable $\tilde{\theta}$ with a probability mass function $g_{n}$ defined on $\Theta_{d}$. Since $g_{n}(\vartheta_{i})>0$ for every $i$, the random variables $(e^{k\tilde{\theta}})_{k\geq 0}$ are linearly independent. Furthermore, $u(\tilde{\theta})$ cannot be constant because $u(\vartheta_{d})=0$ and $\mathbf{u}$ is not the zero vector. Therefore, equation (C.4) can not hold.
We now prove that $\mathbf{V}_{n}$ converge to $\mathbf{V}_{\tilde{g}}$, $P$-a.s. Since the random matrices $\mathbf{V}_{n}$ and $\mathbf{V}_{\tilde{g}}$ have equal finite dimensions, then the convergence follows if we prove that
$\mathbf{V}_{n,i,j}\rightarrow\mathbf{V}_{\tilde{g},i,j}\quad P-a.s.,$
for every $i$ and $j$. Since $\mathbf{V}_{n,i,j}$ is a continuous function of $g_{n}$ and since $g_{n}$ converges to $\tilde{g}$, $P$-a.s, then $\mathbf{V}_{n,i,j}$ converges to $\mathbf{V}_{\tilde{g},i,j}$, $P$-a.s. ∎
Theorem C.2.
If $(b_{n})$ is a sequence of strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1,$ (C.5)
then, as $n\rightarrow+\infty$, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges, in the sense of almost-sure conditional convergence with respect to the filtration $(\mathcal{G}_{n})$, to a centered Gaussian kernel with random covariance matrix $\mathbf{V}_{\tilde{g}}$ as in (C.2).
Proof.
The proof is based on Theorem B.2. Fix a vector $\mathbf{u}$ in $\mathbb{R}^{d-1}$ such that $||\mathbf{u}||=1$. For each $n\geq 0$, define
$\mathcal{F}_{n,0}=\mathcal{G}_{n},\quad\mathcal{F}_{n,j}=\mathcal{G}_{n+j-1}\mbox{ for }j\geq 1,$
$M_{n,0}=0,\quad M_{n,j}=b_{n}^{1/2}(\langle g_{n},\mathbf{u}\rangle-\langle g_{n+j-1},\mathbf{u}\rangle)\mbox{ for }j\geq 1.$
For every $n\geq 0$, the sequence $(M_{n,j})_{j\geq 0}$ is a bounded martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 0}$, converging in $L^{1}$ to
$M_{n,\infty}=b_{n}^{1/2}\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle.$
Let
$X_{n,j}=M_{n,j}-M_{n,j-1}=b_{n}^{1/2}\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle,$
$X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|=b_{n}^{1/2}\sup_{j\geq 1}|\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle|,$
and
$U_{n}=\sum_{j=0}^{\infty}X_{n,j}^{2}=b_{n}\sum_{k\geq n}(\langle\mathbf{g}_{k}-\mathbf{g}_{k-1},\mathbf{u}\rangle)^{2}.$
The thesis follows from Theorem B.2, if we can verify that
(a) the sequence $(X^{*}_{n})$ is dominated in $L^{1}$ and converges to zero $\mathds{P}$-a.s. as $n\rightarrow+\infty$;
(b) $U_{n}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u}$, $P$-a.s. as $n\rightarrow+\infty$.
(a) Notice that $\langle\mathbf{g}_{k+1}-\mathbf{g}_{k},\mathbf{u}\rangle=\alpha_{k}\langle\mathbf{g}_{k-1}(\cdot\mid Y_{k})-\mathbf{g}_{k-1},\mathbf{u}\rangle$ is bounded uniformly with respect to $k$. Thus, $X^{*}_{n}\leq b_{n}^{1/2}\sup_{k\geq n}\alpha_{k}|\langle g_{k-1}(\cdot\mid Y_{k})-g_{k-1},\mathbf{u}\rangle|$ is uniformly bounded. Furthermore, it converges to zero $P$-a.s., since $b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0$ by assumption.
(b) The proof is based on Theorem B.4 (b). Let $Z_{n}=\alpha_{n}^{-2}(\langle g_{n}-g_{n-1},\mathbf{u}\rangle)^{2}$ and $a_{k}=(b_{k}\alpha_{k})^{-2}$. We can write that
$E_{n-1}(Z_{n})$ $=E_{n-1}\left(\alpha_{n}^{-2}(\langle g_{n}-g_{n-1},\mathbf{u}\rangle)^{2}\right)$
$=\sum_{z\in\mathbb{N}_{0}}\sum_{i,j=1}^{d-1}u_{i}u_{j}\left(g_{n}(\vartheta_{i}\mid z)g_{n}(\vartheta_{j}\mid z)-g_{n}(\vartheta_{i})g_{n}(\vartheta_{j})\right)p_{g_{n}}(z)=\mathbf{u}^{T}V_{n}\mathbf{u}.$
Thus, by Lemma C.1,
$E_{n-1}\left(Z_{n}\right)\rightarrow\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
Moreover, $b_{n}\uparrow+\infty$ by assumption, and
$\sum_{k=1}^{\infty}\frac{E(Z_{k}^{2})}{a_{k}^{2}b_{k}^{2}}=\sum_{k=1}^{\infty}\alpha_{k}^{4}b_{k}^{2}E\left(\left[\langle\mathbf{g}_{k-1}(\cdot\mid Y_{k})-\mathbf{g}_{k-1},\mathbf{u}\rangle\right]^{4}\right)<+\infty,$
where $(\mathbf{g}(\cdot\mid Y_{k}))_{i}=g(\vartheta_{i}\mid Y_{k})$. Since, by assumption, $b_{n}\sum_{k\geq n}a_{k}^{-1}b_{k}^{-2}\rightarrow 1$, then by Theorem B.4 (b),
$b_{n}\sum_{k\geq n}(\langle g_{k}-g_{k-1},\mathbf{u}\rangle)^{2}=b_{n}\sum_{k\geq n}\frac{Z_{k}}{a_{k}b_{k}^{2}}\rightarrow\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
It follows by Theorem B.2 that $M_{n,\infty}=b_{n}^{1/2}(\langle g_{n}-\tilde{g},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(g_{n}-\tilde{g})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,V_{\tilde{g}})$. ∎
of Theorem 2.3.
The proof is based on Theorem C.2 and Theorem B.5. We first prove that (C.5) holds under the assumptions of Theorem 2.3. We can write that
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1.$
In particular, $\alpha_{n}^{4}b_{n}^{2}\rightarrow 0$, as $n\rightarrow\infty$. Since $(\alpha_{n})$ is non increasing, then
$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=b_{n}\alpha_{n}^{2}\rightarrow 0.$
Thus, (C.5) holds. Let us write
$\hat{\theta}_{g}(y)=h_{y}(\mathbf{g}),$
where
$h_{y}(\mathbf{g})=(y+1)\frac{p_{g}(y+1)}{p_{g}(y)}=(y+1)\frac{\sum_{i=1}^{d-1}k_{y+1}(\vartheta_{i})\mathbf{g}_{i}+k_{y+1}(\vartheta_{d})(1-\sum_{i=1}^{d-1}\mathbf{g}_{i})}{\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})\mathbf{g}_{i}+k_{y}(\vartheta_{i})(1-\sum_{i=1}^{d-1}\mathbf{g}_{i})}$
is seen as a function of the vector $\mathbf{g}\in\Delta^{d}$. The function $h_{y}$ is differentiable with respect to $\mathbf{g}$, and, for $i=1,\dots,d-1$,
$\frac{\partial h_{y}(\mathbf{g})}{\partial\mathbf{g}_{i}}$ $=(y+1)\frac{k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})}{\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})g(\vartheta_{i})+k_{y}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i}))}$
$-(y+1)\frac{(\sum_{i=1}^{d-1}k_{y+1}(\vartheta_{i})g(\vartheta_{i})+k_{y+1}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i})))(k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d}))}{(\sum_{i=1}^{d-1}k_{y}(\vartheta_{i})g(\vartheta_{i})+k_{y}(\vartheta_{d})(1-\sum_{i=1}^{d-1}g(\vartheta_{i})))^{2}}$
$=\frac{y+1}{p_{g}(y)}\left[k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})-\frac{p_{g}(y+1)}{p_{g}(y)}(k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d}))\right]$
$=\hat{\theta}_{g}(y)\left[\frac{k_{y+1}(\vartheta_{i})-k_{y+1}(\vartheta_{d})}{p_{g}(y+1)}-\frac{k_{y}(\vartheta_{i})-k_{y}(\vartheta_{d})}{p_{g}(y)}\right].$
Let us denote by $\nabla_{g}h_{y}(\mathbf{g})$ the row vector of partial derivatives of $h_{y}$. Since $\Theta_{d}$ is finite, then the partial derivatives are bounded, uniformly with respect to $\mathbf{g}\in\Delta^{d}$. Thus, $h_{y}$ is uniformly differentiable. By Theorem C.2, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to a centered normal kernel $\mathcal{N}(0,\mathbf{V}_{\tilde{\mathbf{g}}})$, with random covariance matrix $\mathbf{V}_{\tilde{\mathbf{g}}}$ as in (C.2). By Theorem B.5, $b_{n}^{1/2}(h_{y}(\mathbf{g}_{n})-h_{y}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence, to a Gaussian kernel, with covariance matrix
$\nabla_{\tilde{\mathbf{g}}}h_{y}(\tilde{\mathbf{g}})\mathbf{V}_{\tilde{g}}\nabla_{\tilde{g}}h_{\tilde{g}}^{T}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{i,j=1}^{d-1}\left[\frac{k(y+1\mid\vartheta_{i})-k(y+1\mid\vartheta_{d})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d})}{p_{\tilde{g}}(y)}\right]$
$\quad\left[\frac{k(y+1\mid\vartheta_{j})]-k(y+1|\vartheta_{d})}{p_{\tilde{g}}(y+1)}-\frac{k(y\mid\vartheta_{j})-k(y\mid\vartheta_{d})}{p_{\tilde{g}}(y)}\right]$
$\left[\sum_{z\in\mathbb{N}_{0}}\tilde{g}(\vartheta_{i}\mid z)\tilde{g}(\vartheta_{j}\mid z)p_{\tilde{g}}(z)-\tilde{g}(\vartheta_{i})\tilde{g}(\vartheta_{j})\right]$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i}\mid z)(\frac{\tilde{g}(\vartheta_{i}\mid y+1)}{\tilde{g}(\vartheta_{i})}-\frac{\tilde{g}(\vartheta_{d}\mid y+1)}{\tilde{g}(\vartheta_{d})}-\frac{\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}+\frac{\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})})\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i})(\frac{\tilde{g}(\vartheta_{i}\mid y+1)}{\tilde{g}(\vartheta_{i})}-\frac{\tilde{g}(\vartheta_{d}\mid y+1)}{\tilde{g}(\vartheta_{d})}-\frac{\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}+\frac{\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})})\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i}\mid z)(\frac{\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})})-(1-\tilde{g}(\vartheta_{d}\mid z))\frac{\tilde{g}(\vartheta_{d}\mid y+1)-\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})}\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d-1}\tilde{g}(\vartheta_{i})\frac{\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)}{\tilde{g}(\vartheta_{i})}-(1-\tilde{g}(\vartheta_{d}))\frac{\tilde{g}(\vartheta_{d}\mid y+1)-\tilde{g}(\vartheta_{d}\mid y)}{\tilde{g}(\vartheta_{d})}\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d}\frac{\tilde{g}(\vartheta_{i}\mid z)}{\tilde{g}(\vartheta_{i})}(\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y))\right]^{2}$
$-\hat{\theta}_{\tilde{g}}(y)^{2}\left[\sum_{i=1}^{d}\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y)\right]^{2}$
$=\hat{\theta}_{\tilde{g}}(y)^{2}\sum_{z\in\mathbb{N}_{0}}p_{\tilde{g}}(z)\left[\sum_{i=1}^{d}\frac{\tilde{g}(\vartheta_{i}\mid z)}{\tilde{g}(\vartheta_{i})}(\tilde{g}(\vartheta_{i}\mid y+1)-\tilde{g}(\vartheta_{i}\mid y))\right]^{2}.$
C.4 Proof of Theorem 2.4
The proof of Theorem 2.4 is based on lemmas C.3, C.7 and on the rate of convergence of $g_{n}$ to $g_{\eta}^{*}$, under i.i.d. observations. This last result is known in the literature (Martin and Tokdar, 2009), but we think that the proof is not complete. We provide for completeness the proof in Lemma C.5.
In this section, we will occasionally represent probability mass functions $g$ on $\Theta_{\eta}$ as vectors $\mathbf{g}$, taking values either in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0\;(i=1,\dots,d_{\eta}-1),\;\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$ or in $\mathbb{R}^{d_{\eta}}$, depending on the context, which will be specified when needed. Similarly, the Poisson kernel is represented as a vector in $\Delta_{\eta}$ or $\mathbb{R}^{d_{\eta}}$, by defining $\mathbf{k}_{y}(\vartheta_{j})=k(y\mid\vartheta_{j})$. We denote component-wise multiplication of vectors by $\circ$.
Lemma C.3.
If
$\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{*}}(y)<+\infty,$ (C.6)
then the $d_{\eta}$-KL-discretization $g_{\eta}^{*}$ of $G^{*}$ satisfies for every $i=1,\dots,d_{\eta}$ the equation in $g$:
$g(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)=0.$ (C.7)
Proof.
Representing probability mass functions on $\Theta_{\eta}$ as vectors in $\Delta_{\eta},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a function on $\Delta_{\eta}$. The assumption (C.6) implies that
$\int_{0}^{\infty}e^{\theta(d_{\eta}^{2}-1)}G^{*}(d\theta)$ $=\int_{0}^{\infty}\sum_{y\in\mathbb{N}_{0}}e^{\log(d_{\eta}^{2})y}e^{-\theta}\frac{\theta^{y}}{y!}G^{*}(d\theta)$ (C.8)
$=\sum_{y\in\mathbb{N}_{0}}e^{2y\log d_{\eta}}p_{G^{*}}(y)<+\infty.$
Thus, for every $i=1,\dots,d_{\eta}$,
$\frac{\partial}{\partial\mathbf{g}_{i}}\ell(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\sum_{y\in\mathbb{N}_{0}}\left(\frac{\vartheta_{d_{\eta}}}{\vartheta_{1}}\right)^{y}\int_{\Theta}\frac{e^{-\theta}\theta^{y}}{y!}{G^{*}}(\mathrm{d}\theta)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}-1)}G^{*}(\mathrm{d}\theta)<+\infty$
by (C.8) Analogously, for every $i=1,\dots,d_{\eta}$,
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}^{2}-1)}G^{*}(\mathrm{d}\theta)<+\infty.$ (C.9)
Thus, $\ell(\mathbf{g})$ is twice continuously differentiable on $\Delta^{d}$. Denoting by $k_{y}(\vartheta_{i})=k(y\mid\vartheta_{i})$ for $i=1,\dots,\vartheta_{d_{\eta}}$, we can write
$\nabla_{\mathbf{g}}KL(p_{G^{*}}||p_{\mathbf{g}})=-\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y).$
The $(i,j)$ element of the Hessian matrix $\mathbf{M}$ of $\ell(\mathbf{g})$ is
$\mathbf{M}_{i,j}$ $=\sum_{y\in\mathbb{N}_{0}}\frac{(k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}}))(k(y\mid\vartheta_{j})-k(y\mid\vartheta_{d_{\eta}}))}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y).$
for $i,j=1,\dots,d-1$.
To prove (C.7), we apply the method of Lagrange multipliers for the constrained minimum of twice continuously differentiable functions. The solution $\mathbf{g}_{\eta}^{*}$ satisfies, for same $\lambda_{i}\geq 0$ ($i=1,\dots,d_{\eta}$) with $\lambda_{i}>0$ for at least one $i$,
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}+\lambda_{d_{\eta}}=0\quad(i=1,\dots,d_{\eta}-1),$ (C.10)
and
$\lambda_{1}g_{\eta}^{*}(\vartheta_{1})=0,\;\dots,\;\lambda_{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{d_{\eta}-1})=0,\;\lambda_{d_{\eta}}(\sum_{i=1}^{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{i})-1)=0.$ (C.11)
Multiplying (C.10) by $g_{\eta}^{*}(\vartheta_{i})$ and taking the sum with respect to $i=1,\dots,d_{\eta}-1$, we obtain by (C.11)
$\lambda_{d_{\eta}}=-\sum_{i=1}^{d_{\eta}-1}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)+(1-g_{\eta}^{*}(\vartheta_{d_{\eta}}))\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y).$
Thus
$\lambda_{d_{\eta}}-\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-1.$
Substituting in (C.10), we obtain
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}-1=0\quad(i=1,\dots,d_{\eta}-1),$
Thus, by (C.11),
$g_{\eta}^{*}(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)-1\right)=0$ (C.12)
for every $i=1,\dots,d_{\eta}-1$. Since
$\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=1,$
then (C.12) holds also for $i=d_{\eta}$.
$\square$
Lemma C.4.
Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$:
$\mathbf{g}_{n+1}=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$ (C.13)
with
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)\right),$
and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d_{\eta}}$, converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}_{G^{*}}$-a.s.
Proof.
Let $J$ be the set of vectors in $\Delta^{d_{\eta}}$ that satisfy (C.7). According to Theorem 2.2.1 in Chen (2005), if the following conditions hold:
  • S1
    $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$;
  • S2
    $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$ and $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$;
  • S3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
  • S4
    $\mathbf{h}$ is measurable and locally bounded,
then
$d(\mathbf{g}_{n},J)\rightarrow 0,$ (C.14)
$\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where $d$ denotes the Euclidean distance in $\mathbb{R}^{d_{\eta}}$.
S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and H4 is obvious.
We now prove S2. Let
$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{G^{*}}(y).$
By (C.6), the function $\mathscr{l}(\mathbf{g})=KL(p_{G^{*}}||p_{\mathbf{g}})$ is twice continuously differentiable and strictly convex in $\Delta^{d_{\eta}}$, as proved in Lemma C.3, with gradient
$\nabla_{\mathbf{g}}\mathscr{l}(\mathbf{g})=-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{g}(y)}p_{G^{*}}(y).$
The set $J$ is closed. We now prove that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$. The set $J$ is made of a finite number of isolated points. Indeed, let $\mathbf{g}$ be a point in $J$ and let $N=\{i\in\{1,\dots,{d_{\eta}}\}:\mathbf{g}_{i}=0\}$. We prove that the conditions in (C.7) identify the values of $\mathbf{g}_{i}$ for $i\not\in N$. The property is obvious if $N$ contains ${d_{\eta}}-1$ points. Otherwise, assume by contradiction that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\mathbf{g}_{i}^{\prime}=\mathbf{g}_{i}^{\prime\prime}=0$ for $i\in N$ and $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i\not\in N$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$. Then, for every $a\in(0,1)$
$1=\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
$<\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
$=1,$
which leads to a contradiction. Thus $J$ is made of a finite number of isolated points. It follows that $\{\mathscr{l}(\mathbf{g}):\mathbf{g}\in J\}$ is nowhere dense in $\mathbb{R}$. To conclude the proof of S2, note that, for $\mathbf{g}\in J^{c}$
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right)$
$=-\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)<0.$
Hence S2 holds, and, by (C.14)
$\inf_{\mathbf{g}\in J}||\mathbf{g}_{n}-\mathbf{g}||\rightarrow 0,$
$\mathds{P}_{G^{*}}$-a.s. as $n\rightarrow\infty$.
We now prove that $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s. to the minimizer $\mathbf{g}_{\eta}^{*}$ of the Kullback-Leibler divergence $KL(p_{G^{*}}||p_{g})$.
First, notice that, since $\mathbf{g}_{n+1}-\mathbf{g}_{n}$ converges to zero $\mathds{P}$-a.s. and since $J$ is made of isolated points, then $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s. to a random vector taking values in $J$.
The minimizer $\mathbf{g}_{\eta}^{*}$ of $\mathscr{l}(\mathbf{g})$ satisfies
$(\mathbf{g}_{\eta}^{*}-\mathbf{g})^{T}\nabla\mathscr{l}(\mathbf{g})<0$
for every $\mathbf{g}\neq\mathbf{g}_{\eta}^{*}$. Hence
$\sum_{i=1}^{{d_{\eta}}-1}(\mathbf{g}_{\eta,i}^{*}-\mathbf{g}_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{g}(y)}p_{G^{*}}(y)>0,$
which implies that
$\sum_{i=1}^{{d_{\eta}}}\mathbf{g}_{i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0,$
or, equivalently
$\sum_{i=1}^{{d_{\eta}}}(\mathbf{g}_{i}^{*}-\mathbf{g}_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0.$ (C.15)
We now show that that, for every $\delta>0$, $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},J_{\delta})=0)=0$, where $J_{\delta}=\{\mathbf{g}\in J:d(\mathbf{g},\mathbf{g}_{\eta}^{*})>\delta\}$.
If $\mathbf{g}\in J_{\delta}$, then there exists $\vartheta_{i}$ such that $g(\vartheta_{i})=0$ and $g_{\eta}^{*}(\vartheta_{i})\neq 0$. By contradiction, suppose that there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),J_{\delta})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d_{\eta}}\log(g_{\eta}^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g_{\eta}^{*}(\vartheta_{i})$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$. Let $\overline{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})+\nabla_{\mathbf{g}}\mathcal{L}(\mathbf{g}_{n})^{T}(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$+(\mathbf{g}_{n+1}-\mathbf{g}_{n})^{T}\nabla^{2}_{\mathbf{g}}\mathcal{L}(\overline{\mathbf{g}}_{n})(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $Y_{1},\dots,Y_{n}$, we can write that
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\mid\mathcal{G}_{n}\right),$
with
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\mid\mathcal{G}_{n}\right)=0,$
and
$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\mid\mathcal{G}_{n}\right)=\liminf_{n}\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})\mathbf{g}_{n,i}p_{G^{*}}(y)}{p_{\mathbf{g}_{n}}(y)}-\mathbf{g}_{n,i}\right)$
$=\liminf_{n}\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)-1\right)$
$>0,$
by (C.15). Thus there exists a constant $c>0$ such that
$E_{G^{*}}\left(-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand,
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{\eta,i}^{*}\frac{\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{p_{\mathbf{g}_{n}}(Y_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\left(\frac{k_{Y_{n+1}}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}k_{Y_{n+1}}(\vartheta_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{y\in\mathbb{N}_{0}}\sum_{i,k}\frac{k_{y}(\vartheta_{i})^{2}}{(\sum_{\vartheta_{k}}k_{y}(\vartheta_{k})g_{n}(\vartheta_{k}))^{2}}\int_{\Theta}k_{y}(\theta)G^{*}(\mathrm{d}\theta)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{g_{n}}(y)^{2}}p_{G^{*}}(y)\right)$
$=o(\alpha_{n}).$
where the last equality holds by (C.6), as proved in Lemma C.3. It follows that,
$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
Lemma C.5.
Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ be the representations of $g_{n}$ ($n\geq 0$) and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$. Under the assumptions of Theorem 2.4 with $\gamma<1$, for every $\delta<2-1/\gamma$
$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
Proof.
The random vectors $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ take values in the compact set $\{\mathbf{v}\in\mathbb{R}^{d_{\eta}}:||\mathbf{v}||\leq 1\}$. By Lemma C.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{d_{\eta}}$ converge $P_{G^{*}}$-a.s. to $\mathbf{g}_{\eta}^{*}\in\Delta^{d_{\eta}}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$ also when seen as vectors in $\mathbb{R}^{d_{\eta}}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{d_{\eta}}$:
$\mathbf{g}_{n+1}$ $=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\bf 1\right)$
$=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$
with initial value $\mathbf{g}_{0}$, where
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
and
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right).$
The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that the following conditions hold:
  • T1
    $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$;
  • T2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}\epsilon_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.;
  • T3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}_{\eta}^{*}$. Let $\mathbf{H}$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}_{\eta}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}_{\eta}^{*})+\mathbf{r}(g),\quad\mathbf{r}(g_{\eta}^{*})=\mathbf{0},\quad||\mathbf{r}(\mathbf{g})||=o(||\mathbf{g}-\mathbf{g}_{\eta}^{*}||).$
    All the eigenvalues of the matrix $\mathbf{H}$ have negative real parts.
Remark C.6.
Theorem 3.1.1 in Chen (2005) also requests the following condition: T4 There exists a continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}_{\eta}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}_{\eta}^{*}$. In our case, this property has been proved in Lemma C.4.
T1 is obvious. To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}\epsilon_{k})_{n\geq 1}$ is a martingale. Moreover, there exists a constant $C$ such that, for every $i=1,\dots,{d_{\eta}}$,
$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i})^{2})$
$\leq C\sum_{k=1}^{\infty}\left(\frac{1}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}_{\eta}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}_{\eta,i}^{*}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y).$
The matrix $\mathbf{H}$ can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}_{\eta}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)$. We now prove that $\mathbf{M}$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{d_{\eta}}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{d_{\eta}}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}(\vartheta_{i})k_{y}(\vartheta_{j})}{p_{g_{\eta}^{*}}(y)^{2}}p_{G^{*}}(y)=\sum_{y\in\mathbb{N}_{0}}\frac{p_{G^{*}{(y)}}}{p_{g_{\eta}^{*}}(y)^{2}}\left(\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{d_{\eta}}\mathbf{u}_{i}k_{y}(\vartheta_{i})=0$ for every $y\in\mathbb{N}_{0}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
Lemma C.7.
Under the assumptions of Theorem 2.4, there exists a constant $C=C(d,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
where $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ are the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$.
Proof.
We can write that
$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}p_{g_{\eta}^{*}}(y)$
$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})}\right)^{2}p_{G^{*}}(y)$
$\leq\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)\right)^{2}}{(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i}))^{2}}p_{G^{*}}(y)$
$+\sum_{y\in\mathbb{N}_{0}}\frac{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g_{\eta}^{*}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)+2\sum_{y\in\mathbb{N}_{0}}\frac{\vartheta_{d_{\eta}}^{2}\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}}{\left(\sum_{i=1}^{d_{\eta}}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{\eta}^{*}(\vartheta_{i})\right)^{2}}p_{G^{*}}(y)$
$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{-\theta}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\left(\frac{\theta\vartheta_{d_{\eta}}^{2}}{\vartheta_{1}^{2}}\right)^{y}G^{*}(d\theta)$
$\leq 4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)$
$\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
with $C=4{d_{\eta}}\vartheta_{d_{\eta}}^{2}e^{2(\vartheta_{d_{\eta}}-\vartheta_{1})}\int_{0}^{\infty}e^{\theta({d_{\eta}}^{2}-1)}G^{*}(d\theta)<+\infty,$ by (C.8). Thus,
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2},$
with $C<+\infty$. ∎
of Theorem 2.4.
Let $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ denote the representation of $g_{n}$ and $g_{\eta}^{*}$ in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$. The sequence $(\mathbf{g}_{n})$ satisfies the stochastic approximation (C.13), with $\mathbf{g}_{0}\in\mathring{(\Delta_{\eta})}$. By Lemma C.4, $\mathbf{g}_{n}$ converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}^{*}$-a.s. It follows that $g_{n}(\vartheta_{j})$ converges to $g_{\eta}^{*}(\vartheta_{j})$ $\mathds{P}^{*}$-a.s. for every $\vartheta_{j}\in\Theta_{\eta}$, which imples that $\hat{\theta}_{g_{n}}(y)$ converges to $\hat{\theta}_{g^{*}_{\eta}}(y)$ for every $y\in\mathbb{N}_{0}$.
Now assume that $\gamma<1$. Denoting by $\mathbf{g}_{n}$ and $\mathbf{g}_{\eta}^{*}$ the representations of $g_{n}$ and $g_{\eta}^{*}$ in $\mathbb{R}^{d_{\eta}}$, we can write, by Lemma C.5 that
$||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}),$
for every $\delta<2-1/\gamma$. By Lemma C.7,
${\rm Regret}(g_{n},g_{\eta}^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}_{\eta}^{*}||^{2}=o(n^{-\delta}).$
$\square$
D Numerical illustrations
D.1 Synthetic data
[…]
Table 3: Weibull prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=454,169$ $d_{\eta}=104,053$ $d_{\eta}=43,755$ $d_{\eta}=23,612$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=11,354.225$ $\vartheta_{d_{\eta}}=52,02.65$ $\vartheta_{d_{\eta}}=3,281.625$ $\vartheta_{d_{\eta}}=2,361.2$
$n=100$ $d_{\eta}=369,148$ $d_{\eta}=84,375$ $d_{\eta}=35,423$ $d_{\eta}=19,090$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=9,228.7$ $\vartheta_{d_{\eta}}=4,218.75$ $\vartheta_{d_{\eta}}=2,656.725$ $\vartheta_{d_{\eta}}=1,909$
$n=200$ $d_{\eta}=450,573$ $d_{\eta}=103,220$ $d_{\eta}=43,402$ $d_{\eta}=23,420$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=11,264.325$ $\vartheta_{d_{\eta}}=5,161$ $\vartheta_{d_{\eta}}=3,255.15$ $\vartheta_{d_{\eta}}=2,342$
$n=500$ $d_{\eta}=431,768$ $d_{\eta}=98,865$ $d_{\eta}=41,558$ $d_{\eta}=22,419$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=10,794.2$ $\vartheta_{d_{\eta}}=49,43.25$ $\vartheta_{d_{\eta}}=31,16.85$ $\vartheta_{d_{\eta}}=2,241.9$
Table 4: Weibull prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 1.572 1.575 1.578 1.581
MAD 1.319 1.322 1.325 1.328
$n=100$
RMSE 1.677 1.72 1.75 1.772
MAD 1.335 1.366 1.386 1.4
$n=200$
RMSE 1.335 1.336 1.336 1.337
MAD 1.064 1.064 1.065 1.065
$n=500$
RMSE 1.348 1.348 1.349 1.349
MAD 1.071 1.071 1.071 1.071
[…]
Table 5: Weibull prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 1.574 1.581 1.582 6.13 14.49
MAD 1.323 1.328 1.329 5.706 12.518
CPU time 1.243 0.24 0.12 0.025 0.013
$n=100$
RMSE 1.667 1.675 1.677 4.496 11.476
MAD 1.335 1.335 1.335 4.101 10.536
CPU time 1.208 0.24 0.12 0.024 0.013
$n=200$
RMSE 1.33 1.336 1.336 6.039 13.984
MAD 1.062 1.064 1.065 5.733 11.949
CPU time 1.231 0.242 0.122 0.025 0.013
$n=500$
RMSE 1.348 1.348 1.348 5.869 12.96
MAD 1.071 1.071 1.072 5.553 11.152
CPU time 1.214 0.241 0.12 0.025 0.013
Table 6 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 3, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization in the second column of Table 5, which also provides CPU time.
Table 6: Weibull prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 3.154 1.542 1.559 1.554 1.581
RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
RMSE   $n=200$ 3.236 1.273 1.248 1.262 1.336
RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
MAD     $n=50$ 2.32 1.262 1.258 1.295 1.328
MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
MAD     $n=200$ 1.917 0.999 1.001 1.014 1.064
MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
[…]
Table 7: Uniform prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=71,674$ $d_{\eta}=15,988$ $d_{\eta}=6,596$ $d_{\eta}=3,504$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,791.85$ $\vartheta_{d_{\eta}}=799.4$ $\vartheta_{d_{\eta}}=494.7$ $\vartheta_{d_{\eta}}=350.4$
$n=100$ $d_{\eta}=67,264$ $d_{\eta}=14,986$ $d_{\eta}=6,177$ $d_{\eta}=3,279$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,681.6$ $\vartheta_{d_{\eta}}=749.3$ $\vartheta_{d_{\eta}}=463.275$ $\vartheta_{d_{\eta}}=327.9$
$n=200$ $d_{\eta}=52,298$ $d_{\eta}=11,593$ $d_{\eta}=4,760$ $d_{\eta}=2,519$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,307.45$ $\vartheta_{d_{\eta}}=579.65$ $\vartheta_{d_{\eta}}=357$ $\vartheta_{d_{\eta}}=251.9$
$n=500$ $d_{\eta}=44,591$ $d_{\eta}=9,850$ $d_{\eta}=4,034$ $d_{\eta}=2,130$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,114.775$ $\vartheta_{d_{\eta}}=492.5$ $\vartheta_{d_{\eta}}=302.55$ $\vartheta_{d_{\eta}}=213$
Table 8: Uniform prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 0.724 0.725 0.727 0.728
MAD 0.615 0.616 0.617 0.618
$n=100$
RMSE 0.798 0.798 0.799 0.8
MAD 0.648 0.65 0.651 0.653
$n=200$
RMSE 0.728 0.729 0.73 0.731
MAD 0.619 0.62 0.621 0.622
$n=500$
RMSE 0.72 0.72 0.721 0.721
MAD 0.59 0.59 0.59 0.59
[…]
Table 9: Uniform prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 0.724 0.724 0.724 0.737 1.587
MAD 0.615 0.615 0.616 0.619 1.397
CPU time 1.233 0.241 0.125 0.026 0.014
$n=100$
RMSE 0.799 0.804 0.812 0.841 1.514
MAD 0.649 0.653 0.66 0.678 1.302
CPU time 1.310 0.241 0.120 0.026 0.014
$n=200$
RMSE 0.728 0.73 0.731 0.751 1.056
MAD 0.619 0.619 0.619 0.623 0.831
CPU time 1.252 0.239 0.131 0.025 0.013
$n=500$
RMSE 0.72 0.719 0.719 0.721 0.854
MAD 0.59 0.612 0.612 0.64 0.665
CPU time 1.207 0.241 0.119 0.025 0.013
Table 10 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 7, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization considered in the second column of Table 9, which also provides the CPU time. Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
Table 10: Uniform prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 1.354 0.727 0.702 0.703 0.724
RMSE   $n=100$ 1.253 0.818 0.841 0.763 0.804
RMSE   $n=200$ 0.88 0.718 0.729 0.726 0.73
RMSE   $n=500$ 0.88 0.725 0.746 0.72 0.719
MAD     $n=50$ 0.996 0.608 0.562 0.581 0.615
MAD     $n=100$ 0.998 0.673 0.685 0.628 0.653
MAD     $n=200$ 0.742 0.613 0.605 0.623 0.619
MAD     $n=500$ 0.719 0.606 0.633 0.607 0.612
Figure 2: Data $Y_{i}$’s (Uniform prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
[…]
Table 11: Half-Gaussian prior: grid $\Theta_{\eta}$
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=29,100$ $d_{\eta}=6,362$ $d_{\eta}=2,585$ $d_{\eta}=1,355$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=727.5$ $\vartheta_{d_{\eta}}=318.1$ $\vartheta_{d_{\eta}}=193.875$ $\vartheta_{d_{\eta}}=135.5$
$n=100$ $d_{\eta}=19,629$ $d_{\eta}=4,243$ $d_{\eta}=1,709$ $d_{\eta}=889$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=490.725$ $\vartheta_{d_{\eta}}=212.15$ $\vartheta_{d_{\eta}}=128.175$ $\vartheta_{d_{\eta}}=88.9$
$n=200$ $d_{\eta}=16,878$ $d_{\eta}=3,631$ $d_{\eta}=1,456$ $d_{\eta}=755$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=421.95$ $\vartheta_{d_{\eta}}=181.55$ $\vartheta_{d_{\eta}}=109.2$ $\vartheta_{d_{\eta}}=75.5$
$n=500$ $d_{\eta}=17,307$ $d_{\eta}=3,727$ $d_{\eta}=1,496$ $d_{\eta}=776$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=432.675$ $\vartheta_{d_{\eta}}=186.35$ $\vartheta_{d_{\eta}}=112.2$ $\vartheta_{d_{\eta}}=77.6$
Table 12: Half-Gaussian prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 0.569 0.57 0.571 0.572
MAD 0.462 0.463 0.464 0.465
$n=100$
RMSE 0.633 0.633 0.627 0.624
MAD 0.532 0.535 0.538 0.541
$n=200$
RMSE 0.471 0.471 0.472 0.472
MAD 0.372 0.373 0.374 0.375
$n=500$
RMSE 0.536 0.537 0.537 0.538
MAD 0.425 0.425 0.426 0.427
[…]
Table 13: Half-Gaussian prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 0.569 0.57 0.57 0.58 0.632
MAD 0.461 0.462 0.462 0.463 0.489
CPU time 1.234 0.242 0.134 0.027 0.013
$n=100$
RMSE 0.633 0.633 0.633 0.635 0.687
MAD 0.541 0.541 0.542 0.543 0.594
CPU time 1.226 0.241 0.12 0.024 0.013
$n=200$
RMSE 0.471 0.471 0.471 0.476 0.486
MAD 0.372 0.372 0.372 0.375 0.385
CPU time 1.208 0.241 0.118 0.024 0.013
$n=500$
RMSE 0.535 0.535 0.535 0.536 0.536
MAD 0.419 0.42 0.424 0.424 0.425
CPU time 1.223 0.241 0.12 0.025 0.013
Table 14 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 11, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization considered in the second column of Table 13, which also provides the CPU time. Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
Table 14: Half-Gaussian prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 1.129 0.623 0.643 0.578 0.57
RMSE   $n=100$ 0.679 0.511 0.484 0.493 0.633
RMSE   $n=200$ 0.611 0.477 0.491 0.483 0.471
RMSE   $n=500$ 0.838 0.555 0.564 0.54 0.535
MAD     $n=50$ 0.773 0.49 0.493 0.467 0.462
MAD     $n=100$ 0.512 0.428 0.391 0.411 0.541
MAD     $n=200$ 0.467 0.374 0.403 0.388 0.372
MAD     $n=500$ 0.508 0.429 0.44 0.43 0.42
Figure 3: ata $Y_{i}$’s (half-Gaussian prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
D.2 Real data: Twitter dataset
We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/. It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts. For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting.
For each tweet, we consider the number of retweets generated within $30$ seconds from the tweet post time. From the first two rows of Table 15, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Table 15 and Figure 4 display the estimates of the number of retweets expected for a tweet.
For each tweet, we focus on the number of followers generated by retweets within $30$ seconds from the tweet post time, namely the total number of followers of fast retweets of a tweet, with the proviso that 00 retweets corresponds to 00 followers generated. From the first two rows of Table 16, $40,287$ tweets have generated 00 followers in $30$ seconds, $67$ tweets have generated $1$ follower in $30$ seconds, etc. Table 16 and Figure 5 display the estimates of the number of followers expected to be generated by the retweets of a tweet.
Table 15: Counts $n_{y}$ of the number of tweets that have $y$ retweets within $30$ seconds from the tweet post time, for $y=0,1,\ldots,77$, and corresponding estimates of the number of retweets expected.
$y$ 0 1 2 3 4 5 6 7
$n_{y}$ 40,259 28,339 21,581 16,479 12,130 9,238 7,193 5,464
NP-EB 0.70 1.52 2.29 2.94 3.81 4.67 5.32 6.32
NP-ML 0.39 2.13 2.93 3.70 4.46 5.33 6.33 7.42
NP-MD 0.75 2.20 3.14 4.15 4.99 5.69 6.39 7.16
P-EB 0.45 1.34 2.23 3.12 4.01 4.90 5.79 6.68
QB-EB 0.61 0.82 1.63 3.08 4.01 4.52 5.02 5.71
$y$ 8 11 14 17 20 23 26 29
$n_{y}$ 4,319 2,291 1,352 817 475 293 188 129
NP-EB 7.18 10.10 11.77 15.33 18.66 21.13 22.40 28.37
NP-ML 8.54 11.79 15.01 18.30 20.89 22.96 25.87 30.47
NP-MD 8.04 10.87 13.02 14.83 16.68 18.82 22.50 29.21
P-EB 7.57 10.24 12.91 15.58 18.25 20.93 23.60 26.27
QB-EB 6.67 10.37 12.69 14.64 17.55 21.38 24.78 27.53
$y$ 32 35 38 41 44 47 50 53
$n_{y}$ 89 69 60 61 36 36 32 21
NP-EB 30.78 32.35 30.55 26.85 48.75 45.33 43.03 41.14
NP-ML 34.88 37.75 39.81 41.73 43.79 46.18 49.13 52.72
NP-MD 35.48 39.08 41.63 44.05 46.58 49.21 51.93 54.70
P-EB 28.94 31.61 34.28 36.96 39.63 42.30 44.97 47.64
QB-EB 30.26 33.32 36.59 39.71 42.50 45.05 47.54 50.09
$y$ 56 59 62 65 68 71 74 77
$n_{y}$ 24 16 14 13 14 11 2 5
NP-EB 42.75 48.75 45.00 91.39 14.79 19.64 225.00 62.40
NP-ML 56.58 60.08 62.88 65.06 66.84 68.38 69.77 71.01
NP-MD 57.47 60.21 62.84 65.31 67.54 69.49 71.15 72.53
P-EB 50.31 52.99 55.66 58.33 61.00 63.67 66.35 69.02
QB-EB 52.74 55.40 57.94 60.26 62.33 64.18 65.87 67.45
Figure 4: Estimates of the number of retweets expected for a tweet with $y$ retweets (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 16: Counts $n_{y}$ of the number of tweets that have generated $y$ followers within $30$ seconds from the tweet post time, for $y=0,1,\ldots,450$, and corresponding estimates of the number of followers expected.
$y$ 0 30 60 90 120 150 180 210
$n_{y}$ 40,287 67 87 83 98 82 84 63
NP-EB 0 31.46 55.39 104.16 101.25 147.32 178.85 247.84
NP-ML 0 31.67 61.10 91.01 120.99 149.46 180.19 210.79
NP-MD 0.78 29.93 60.62 93.26 121.19 148.30 181.52 213.79
P-EB 0.56 30.410 60.26 90.12 119.97 149.82 179.67 209.52
QB-EB 0.54 31.41 60.90 91.19 120.68 150.89 180.31 210.85
$y$ 240 270 300 330 360 390 420 450
$n_{y}$ 85 76 65 81 81 63 75 71
NP-EB 275.02 274.57 342.68 314.65 352.09 378.59 370.48 330.31
NP-ML 240.81 270.90 300.96 330.97 361.00 391.05 421.04 450.46
NP-MD 244.11 270.94 300.95 330.98 360.98 390.96 420.81 449.93
P-EB 239.38 269.23 299.08 328.93 358.79 388.64 418.49 448.34
QB-EB 240.40 271.03 300.57 330.71 360.28 390.42 418.99 442.98
Figure 5: Estimates of the number of followers expected to be generated by a tweet whose retweets have $y$ followers (in 30 seconds): NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
D.3 Real data: auto accident dataset
We consider the auto accident data of Efron and Hastie (2021, Table 6.1). This is a benchmark dataset in empirical Bayes. These are one year of claims data for a European automobile insurance company on $9461$ automobile insurance policy holders. From the first two rows of Table 17, $7840$ of the $9461$ policy holders made no claims during the year, $1317$ made a single claim, $239$ made two claims each, etc., continuing to the one person who made seven claims. Interest is in estimating the number of claims expected in a succeeding year by each automobile insurance policy holder. Table 17 displays the estimates of the number of claims expected.
Table 17: Counts $n_{y}$ of number of claims $y$ made in a single year, for $y=0,1,\ldots,7$, and corresponding estimates of the number of claims expected.
Claims $y$ 0 1 2 3 4 5 6 7
Counts $n_{y}$ 7,840 1,317 239 42 14 4 4 1
NP-EB 0.17 0.36 0.53 1.33 1.43 6.00 1.75 0
NP-ML 0.33 1.39 2.31 3.39 4.30 4.93 5.362 5.69
NP-MD 0.15 0.41 0.82 1.00 1.17 1.79 3.24 4.46
P-FB 0.16 0.42 0.68 0.94 1.20 1.47 1.75 2.03
QB-EB 0.12 0.25 0.47 0.91 1.85 3.17 4.33 5.22
E An extension to the multidimensional Poisson compound decision problem
The results in Section 2.2 extend, with minor modifications, to the $k$-dimension coordinate-wise independent Poisson model. We assume that, $\mathbf{Y}=(\mathbf{Y}_{1},\dots,\mathbf{Y}_{k})$ is a $k$-dimensional random vector with independent coordinates $\mathbf{Y}_{i}\sim{\rm Poisson}(\bm{\theta}_{i})$, and denote
$k(\mathbf{y}\mid\bm{\theta})=\prod_{j=1}^{k}e^{-\bm{\theta}_{j}}\frac{\bm{\theta}_{j}^{\mathbf{y}_{j}}}{\mathbf{y}_{j}!},\quad\mathbf{y}=(\mathbf{y}_{1},\dots,\mathbf{y}_{k})\in\mathbb{N}_{0}^{k},\bm{\theta}_{j}\in\Theta_{d}\;(j=1,\dots,k),$ (E.16)
with $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$. Denote $\bm{\theta}=(\bm{\theta}_{1},\dots,\bm{\theta}_{k})\in\Theta_{d}^{k}$ and $\Theta_{d}^{k}=\{{\bm{\vartheta}}_{1},\dots,{\bm{\vartheta}}_{D}\}$, $D=d^{k}$. If $\bm{\theta}$ has probability mass function $g$ on $\Theta_{d}^{k}$, then the marginal probability mass function of $\mathbf{Y}=(\mathbf{Y}_{1},\dots,\mathbf{Y}_{k})$ is
$p_{g}(\mathbf{y})=\sum_{i=1}^{D}k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g({\bm{\vartheta}}_{i}).$ (E.17)
Given a sequence $(\mathbf{Y}_{n})_{n\geq 1}$ of random vectors with independent coordinates $\mathbf{Y}_{n,j}\sim{\rm Poisson}(\bm{\theta}_{n,j})$, we consider the problem of estimating the unknown parameters $\bm{\theta}_{n,j}$.
If we knew that the random vectors $(\bm{\theta}_{n})_{n\geq 1}$ were independent and identically distributed with a known probability mass function $g$, we could estimate $\bm{\theta}_{n,j}$, given that $\mathbf{Y}_{n}=\mathbf{y}$, with
$\hat{\bm{\theta}}_{g,j}(\mathbf{y})=E_{g}[\bm{\theta}_{n,j}|\mathbf{Y}_{n}=\mathbf{y}]=\frac{1}{p_{g}(\mathbf{y})}\sum_{i=1}^{D}e^{-{\bm{\vartheta}}_{i,j}}\frac{{\bm{\vartheta}}_{i,j}^{\mathbf{y}_{j}+1}}{\mathbf{y}_{j}!}\prod_{j^{\prime}\neq j}e^{-{\bm{\vartheta}}_{i,j^{\prime}}}\frac{{\bm{\vartheta}}_{i,j^{\prime}}^{\mathbf{y}_{j^{\prime}}}}{\mathbf{y}_{j^{\prime}}!}=(\mathbf{y}_{j}+1)\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})},$ (E.18)
where $\mathbf{e}_{j}$ is a $k$-dimensional vector with $\mathbf{e}_{j,i}=0$ for $i\neq j$ and $\mathbf{e}_{j,j}=1$.
Being $g$ unknown, we can estimate it recursively as
$g_{n+1}({\bm{\vartheta}}_{i})=(1-\alpha_{n+1})g_{n}({\bm{\vartheta}}_{i})+\alpha_{n+1}\frac{k(\mathbf{Y}_{n+1}\mid{\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i})}{\sum_{i^{\prime}=1}^{D}k(\mathbf{Y}_{n+1}\mid{\bm{\vartheta}}_{i^{\prime}})g_{n}({\bm{\vartheta}}_{i^{\prime}})}\qquad i=1,\dots,D,$ (E.19)
with $g_{0}({\bm{\vartheta}}_{i})>0$ for every $i=1,\dots,D$, and adopt for $\bm{\theta}_{g,j}(\mathbf{y})$ the plug-in estimator
$\hat{\bm{\theta}}_{g_{n},j}(\mathbf{y})=(\mathbf{y}_{j}+1)\frac{p_{g_{n}}(\mathbf{y}+\mathbf{e}_{j})}{p_{g_{n}}(\mathbf{y})}\qquad\mathbf{y}\in\mathbb{N}_{0}^{k},$ (E.20)
where
$p_{g_{n}}(\mathbf{y})=\sum_{i=1}^{D}k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}).$ (E.21)
We now model the sequence of observations $(\mathbf{Y}_{n})$, assuming that, given the information available at time $n$, $\mathbf{Y}_{n+1}$ follows the probability mass function $p_{g_{n}}$. In other words, denoting by $(\mathcal{G}_{n})_{n\geq 0}$ the natural filtration of $(\mathbf{Y}_{n})_{n\geq 1}$, we assume that $\mathbf{Y}_{n+1}\mid\mathcal{G}_{n}\sim p_{g_{n}}$. Under this assumption, the sequence $(g_{n})_{n\geq 0}$ is a martingale with respect to $(\mathcal{G}_{n})_{n\geq 0}$, converging $\mathds{P}$ a.s. as $n\rightarrow\infty$ to a random probability mass function $\tilde{g}$ on $\Theta_{d}^{k}$. In particular, for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$, as $n\rightarrow+\infty$
$\hat{\bm{\theta}}_{g_{n},j}(\mathbf{y})\rightarrow\hat{\bm{\theta}}_{\tilde{g},j}(\mathbf{y})=(\mathbf{y}_{j}+1)\frac{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j})}{p_{\tilde{g}}(\mathbf{y})}\qquad\mathds{P}\,\text{-}\,a.s.$ (E.22)
The next result extends Theorem 2.3 in two directions: the observations are now random vectors, and the sequence of learning rates $(\alpha_{n})$ is no longer required to be increasing.
Theorem E.1.
If $(b_{n})_{n\geq 1}$ are strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{k\geq n}\alpha_{k}^{2}\rightarrow 0,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1,$
then, for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$, $b_{n}^{1/2}(\hat{\bm{\theta}}_{g_{n}}(\mathbf{y})-\hat{\bm{\theta}}_{\tilde{g}}(\mathbf{y}))$ converges, in the sense of almost-sure conditional convergence, as $n\rightarrow+\infty$, to a centered (i.e., zero-mean) Gaussian kernel with the random covariance matrix $\mathbf{W}(\mathbf{y})$, whose $(j,j^{\prime})$ element is given by
$\mathbf{W}_{\tilde{g},j,j^{\prime}}(\mathbf{y})=\hat{\bm{\theta}}_{\tilde{g},j}(\mathbf{y})\hat{\bm{\theta}}_{\tilde{g},j^{\prime}}(\mathbf{y})$ (E.23)
$E_{Z\sim p_{\tilde{g}}}\Bigl{[}\Bigl{(}\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid Z)}{\tilde{g}({\bm{\vartheta}}_{i})}\bigl{(}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\bigr{)}\Bigr{)}\Bigl{(}\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid Z)}{\tilde{g}({\bm{\vartheta}}_{i})}\bigl{(}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\bigr{)}\Bigr{)}\Bigr{]},$ (E.24)
See Supplementary Material F for the proof of Theorem 2.3. For $\alpha_{n}=(\alpha+n)^{-\gamma}$ with $\gamma\in(1/2,1]$, the theorem holds with $b_{n}=(2\gamma-1)/n^{2\gamma-1}$. In particular, with $\alpha_{n}=(1+n)^{-1}$, $b_{n}=n$. Theorem 2.3 provides a tool to obtain asymptotic credible regions for $\hat{\bm{\theta}}_{g}(\mathbf{y})$, for $y\in\mathbb{N}_{0}^{k}$.
We now consider the Bayesian regret under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$. By the same techniques used in the univariate case, the frequentist guarantees can be studied in this more general setting. We give a complete treatment under the simplifying assumption that the “true” probability distribution $G^{*}$ has support $\Theta_{d}^{k}$. Denoting by $g^{*}$ its probability mass function on $\Theta_{d}^{k}$, the regret becomes
$\text{Regret}(g_{n},g^{*})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}||\hat{\bm{\theta}}_{g_{n}}(\mathbf{y})-\hat{\bm{\theta}}_{g^{*}}(\mathbf{y})||^{2}p_{g^{*}}(\mathbf{y}).$ (E.25)
The asymptotic behavior of regret for large $n$ in the multidimensional case follows the same rate of convergence as in the univariate case.
Theorem E.2.
Let $g_{n}$ be as in (E.19) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}_{G^{*}}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$
See Supplementary Material F for the proof of Theorem E.2.
F Proofs for Section E
F.1 Notation
We denote by $\Delta^{D}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{D-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{D-1}\mathbf{v}_{i}\leq 1\}$ the $D$-dimensional simplex, and treat probability mass functions on $\Theta_{d}^{k}$ as vectors in $\Delta^{D}$ or in $\mathbb{R}^{D}$, by setting $\mathbf{g}_{i}=g({\bm{\vartheta}}_{i})$. Analogously, we write $\mathbf{k}_{\mathbf{y},i}=k_{\mathbf{y}}({\bm{\vartheta}}_{i})=k(\mathbf{y}\mid{\bm{\vartheta}}_{i})$ for $i=1,\dots,D$. The probability mass function $p_{g}$ is denoted as $p_{\mathbf{g}}$, when $g$ is considered as a vector. For every $g$, we denote
$g({\bm{\vartheta}}_{i}\mid\mathbf{y})=\frac{k(\mathbf{y}\mid{\bm{\vartheta}}_{i})g({\bm{\vartheta}}_{i})}{p_{g}(\mathbf{y})},$
for $\mathbf{y}\in\mathbb{N}_{0}^{k}$, $i=1,\dots,D$.
F.2 Proof of Theorem E.1
Let us define the $(D-1)\times(D-1)$ random matrices $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$, with $(i,i^{\prime})$ elements (for $i,i^{\prime}=1,\dots,D-1$) given by
$\mathbf{V}_{\tilde{g},i,i^{\prime}}=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{\tilde{g}}(\mathbf{z})-\tilde{g}({\bm{\vartheta}}_{i})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}),$ (F.26)
and
$\mathbf{V}_{n,i,i^{\prime}}=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{g_{n}}(\mathbf{z})-g_{n}({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i^{\prime}}).$ (F.27)
Lemma F.1.
Let $\mathbf{V}_{\tilde{g}}$ and $\mathbf{V}_{n}$ be defined as in (F.26) and (F.27). Then, $\mathds{P}$-a.s, $(V_{n})$ is a sequence of uniformly bounded, positive definite random matrices, converging to $\mathbf{V}_{\tilde{g}}$ as $n\rightarrow\infty$.
Proof.
Since, for every $i,{i^{\prime}}$
$|\mathbf{V}_{n,i,{i^{\prime}}}|\leq 1,$
then the sequence $(\mathbf{V}_{n})$ is uniformly bounded. Furthermore, for every random vector $\mathbf{u}\in\mathbb{R}^{D-1}$ with $||\mathbf{u}||=1$,
$\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}$ $=\sum_{i,{i^{\prime}}=1}^{D-1}\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))\mathbf{u}_{{i^{\prime}}}(g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})-g({\bm{\vartheta}}_{i^{\prime}}))p_{g_{n}}(\mathbf{z})$
$=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\left(\sum_{i=1}^{D-1}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))\right)^{2}p_{g_{n}}(\mathbf{z})\geq 0,$
$\mathds{P}$-a.s. Hence, $\mathds{P}$-a.s., $\mathbf{V}_{n}$ is positive semi-definite. We now prove that it is $\mathds{P}$-a.s. positive definite. By contradiction, suppose that there exists a random vector $\mathbf{u}$ such that for $\omega$ in a set $N^{c}$ with positive probability $\mathbf{u}(\omega)\neq\mathbf{0}$ and $\mathbf{u}^{T}(\omega)\mathbf{V}_{n}(\omega)\mathbf{u}(\omega)=0$. Fix $\omega\in N^{c}$ and, without loss of generality, assume that $p_{g_{n}(\omega)}(\mathbf{z})>0$ for every $\mathbf{z}\in\mathbb{N}_{0}^{k}$. Then, we can write (omitting $\omega$ for simplicity)
$\sum_{i=1}^{D-1}\mathbf{u}_{i}(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i}))=0,$
for every $\mathbf{z}\in\mathbb{N}_{0}^{k}$. Define $u({\bm{\vartheta}}_{i})=\mathbf{u}_{i}$ for $i=1,\dots,D-1$ and $u({\bm{\vartheta}}_{D})=0$. Then, for every $\mathbf{z}=(\mathbf{z}_{1},\dots,\mathbf{z}_{k})\in\mathbb{N}_{0}^{k}$,
$\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{\mathbf{z}_{j}}g_{n}({\bm{\vartheta}}_{i})=\sum_{i=1}^{D}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{\mathbf{z}_{j}}g_{n}({\bm{\vartheta}}_{i})\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}),$
which implies that, for every $\mathbf{x}\in\mathbb{N}_{0}^{k}$,
$\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})e^{\langle\mathbf{x},{\bm{\vartheta}}_{i}\rangle}g_{n}({\bm{\vartheta}}_{i})=\sum_{i=1}^{D}e^{\langle\mathbf{x},{\bm{\vartheta}}_{i}\rangle}g_{n}({\bm{\vartheta}}_{i})\sum_{i=1}^{D}u({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i}).$ (F.28)
To demonstrate that this leads to a contradiction, consider a random vector $\tilde{\bm{\theta}}$ with a probability mass function $g_{n}$ defined on $\Theta_{d}^{k}$. Since $g_{n}({\bm{\vartheta}}_{i})>0$ for every $i$, the random variables $(e^{\langle\mathbf{x},\tilde{\bm{\theta}}_{n}\rangle})_{\mathbf{x}\in\mathbb{N}_{0}^{k}}$ are linearly independent. Furthermore, $u(\tilde{\bm{\theta}})$ cannot be constant because $u({\bm{\vartheta}}_{D})=0$ and $\mathbf{u}$ is not the zero vector. Therefore, equation (F.28) can not hold.
We now prove that $\mathbf{V}_{n}$ converge to $\mathbf{V}$, $P$-a.s. Since the matrices $\mathbf{V}_{n}$ and $\mathbf{V}_{\tilde{g}}$ have equal finite dimensions, then the convergence holds if and only if for every $i$ and $i^{\prime}$
$\mathbf{V}_{n,i,i^{\prime}}\rightarrow\mathbf{V}_{\tilde{g},i,i^{\prime}}\quad\mathds{P}-a.s.,$
which is true since $g_{n}$ converges to $\tilde{g}$ and $\mathbf{V}_{n,i,i^{\prime}}$ is a continuous function of $g_{n}$. ∎
Theorem F.2.
Let $(\mathbf{g}_{n})$ and $\tilde{\mathbf{g}}$ be the representation of $(g_{n})$ and of $\tilde{g}$ in $\Delta^{D}$, respectively. If $(b_{n})$ is a sequence of strictly positive numbers with
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sup_{m\geq n}\alpha_{m}^{2}\rightarrow 0,\quad b_{n}\sum_{m\geq n}\alpha_{m}^{2}\rightarrow 1,$ (F.29)
then, as $n\rightarrow+\infty$, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges, in the sense of almost-sure conditional convergence, to a centered Gaussian kernel with random covariance matrix $\mathbf{V}_{\tilde{g}}$ with $(i,i^{\prime})$ element $\mathbf{V}_{\tilde{g},i,i^{\prime}}$ as in (F.26).
Proof.
The proof is based on Theorem B.2. Fix a vector $\mathbf{u}$ in $\mathbb{R}^{D-1}$ such that $||\mathbf{u}||=1$. For each $n\geq 0$, define
$\mathcal{F}_{n,0}=\mathcal{G}_{n},\quad\mathcal{F}_{n,j}=\mathcal{G}_{n+j-1}\mbox{ for }j\geq 1,$
where $(\mathcal{G}_{n})_{n\geq 0}$ is the natural filtration of $(\mathbf{Y}_{n})_{n\geq 1}$,
$M_{n,0}=0,\quad M_{n,j}=b_{n}^{1/2}(\langle\mathbf{g}_{n},\mathbf{u}\rangle-\langle\mathbf{g}_{n+j-1},\mathbf{u}\rangle)\mbox{ for }j\geq 1.$
For every $n\geq 0$, the sequence $(M_{n,j})_{j\geq 0}$ is a bounded martingale with respect to the filtration $(\mathcal{F}_{n,j})_{j\geq 0}$, converging in $L^{1}$ to
$M_{n,\infty}=b_{n}^{1/2}\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle.$
Let
$X_{n,j}=M_{n,j}-M_{n,j-1}=b_{n}^{1/2}\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle,$
$X_{n}^{*}=\sup_{j\geq 1}|X_{n,j}|=b_{n}^{1/2}\sup_{j\geq 1}|\langle\mathbf{g}_{n+j-1}-\mathbf{g}_{n+j},\mathbf{u}\rangle|,$
and
$U_{n}=\sum_{j=0}^{\infty}X_{n,j}^{2}=b_{n}\sum_{m\geq n}(\langle\mathbf{g}_{m}-\mathbf{g}_{m-1},\mathbf{u}\rangle)^{2}.$
The thesis follows from Theorem B.2, if we can verify that
  • (a)
    the sequence $(X^{*}_{n})$ is dominated in $L^{1}$ and converges to zero $\mathds{P}$-a.s. as $n\rightarrow+\infty$.
  • (b)
    $U_{n}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u}$, $P$-a.s. as $n\rightarrow+\infty$.
(a) Notice that $\langle\mathbf{g}_{m+1}-\mathbf{g}_{m},\mathbf{u}\rangle=\alpha_{m}\sum_{i=1}^{D-1}(g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i}))\mathbf{u}_{i}$ is bounded. Thus, $X^{*}_{n}\leq b_{n}^{1/2}\sup_{m\geq n}\alpha_{m}|\sum_{i=1}^{D-1}(\langle g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i})\mathbf{u}_{i}|$ is uniformly bounded and converges to zero $\mathds{P}$-a.s., since $b_{n}\sup_{m\geq n}\alpha_{m}^{2}\rightarrow 0$ by assumption.
(b) The proof is based on Theorem B.4 (b). Let $Z_{n}=\alpha_{n}^{-2}(\langle\mathbf{g}_{n}-\mathbf{g}_{n-1},\mathbf{u}\rangle)^{2},$ and $a_{n}=(b_{n}\alpha_{n})^{-2}$. We can write that
$E_{n-1}(Z_{n})$ $=E_{n-1}\left(\alpha_{n}^{-2}(\langle\mathbf{g}_{n}-\mathbf{g}_{n-1},\mathbf{u}\rangle)^{2}\right)$
$=\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}}=1}^{D-1}\mathbf{u}_{i}\mathbf{u}_{i^{\prime}}\left(g_{n}({\bm{\vartheta}}_{i}\mid\mathbf{z})g_{n}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})-g_{n}({\bm{\vartheta}}_{i})g_{n}({\bm{\vartheta}}_{i^{\prime}})\right)p_{g_{n}}(\mathbf{z})=\mathbf{u}^{T}\mathbf{V}_{n}\mathbf{u}.$
Thus, by Lemma F.1,
$E_{n-1}\left(Z_{n}\right)\rightarrow\mathbf{\mathbf{u}}^{T}\mathbf{V}_{\tilde{g}}\mathbf{\mathbf{u}},\quad\mathds{P}-a.s.$
Moreover, $b_{n}\uparrow+\infty$ by assumption, and
$\sum_{m=1}^{\infty}\frac{E(Z_{m}^{2})}{a_{m}^{2}b_{m}^{2}}=\sum_{m=1}^{\infty}\alpha_{m}^{4}b_{m}^{2}E\left(\left[\sum_{i=1}^{D-1}g_{m-1}({\bm{\vartheta}}_{i}\mid\mathbf{Y}_{m})-g_{m-1}({\bm{\vartheta}}_{i})\mathbf{u}_{i}\right]^{4}\right)<+\infty.$
Since, by assumption, $b_{n}\sum_{m\geq n}a_{m}^{-1}b_{m}^{-2}\rightarrow 1$, then by Theorem B.4 (b),
$b_{n}\sum_{m\geq n}(\langle\mathbf{g}_{m}-\mathbf{g}_{m-1},\mathbf{u}\rangle)^{2}=b_{n}\sum_{m\geq n}\frac{Z_{m}}{a_{m}b_{m}^{2}}\rightarrow\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u},\quad\mathds{P}-a.s.$
It follows by Theorem B.2 that, as $n\rightarrow\infty$, $M_{n,\infty}=b_{n}^{1/2}(\langle\mathbf{g}_{n}-\tilde{\mathbf{g}},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}\mathbf{V}_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,\mathbf{V}_{\tilde{g}})$.
Now suppose that $\alpha_{n}=(\alpha+n)^{-\gamma}$. With $b_{n}=(2\gamma-1)n^{2\gamma-1}$, the conditions (F.29) hold. Indeed, $b_{n}\uparrow+\infty$,
$\sum_{n=1}^{\infty}\alpha_{n}^{4}b_{n}^{2}=\sum_{n=1}^{\infty}\left(\frac{1}{\alpha+n}\right)^{4\gamma}(2\gamma-1)^{2}n^{4\gamma-2}<+\infty,$
$b_{n}\sup_{m\geq n}\alpha_{m}^{2}=(2\gamma-1)n^{2\gamma-1}\left(\frac{1}{\alpha+n}\right)^{2\gamma}\rightarrow 0,$
and
$b_{n}\sum_{m\geq n}{\alpha_{m}^{2}}=(2\gamma-1)n^{2\gamma-1}\sum_{m\geq n}\left(\frac{1}{\alpha+m}\right)^{2\gamma}\rightarrow 1.$
of Theorem E.1.
The proof is based on Theorem F.2 and Theorem B.5. We treat $\hat{\bm{\theta}}_{g}(\mathbf{y})$ as a vector-valued function of $\mathbf{g}\in\Delta^{D}$, where, for $j=1,\dots,k$, the $j$th component is given by
$\hat{\theta}_{g,j}(\mathbf{y})=\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})=(\mathbf{y}_{j}+1)\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})}=(\mathbf{y}_{j}+1)\frac{\sum_{i=1}^{D-1}k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}{\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}.$
The function $\mathbf{h}_{\mathbf{y}}=[\mathbf{h}^{(1)}_{\mathbf{y}}(\mathbf{g}),\dots,\mathbf{h}^{(k)}_{\mathbf{y}}(\mathbf{g})]^{T}$ is differentiable with respect to $\mathbf{g}$, and, for $i=1,\dots,D-1$,
$\frac{\partial\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})}{\partial\mathbf{g}_{i}}$ $=(\mathbf{y}_{j}+1)\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}[D](1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i}))}$
$-(\mathbf{y}_{j}+1)\frac{(\sum_{i=1}^{D-1}k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i})))(k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D}))}{(\sum_{i=1}^{D-1}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}({\bm{\vartheta}}_{i})+k_{\mathbf{y}}({\bm{\vartheta}}_{D})(1-\sum_{i=1}^{D-1}\mathbf{g}({\bm{\vartheta}}_{i})))^{2}}$
$=\frac{\mathbf{y}_{j}+1}{p_{g}(\mathbf{y})}\left[k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})-\frac{p_{g}(\mathbf{y}+\mathbf{e}_{j})}{p_{g}(\mathbf{y})}(k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D}))\right]$
$=\hat{\theta}_{g,j}(\mathbf{y})\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{p_{g}(\mathbf{y}+\mathbf{e}_{j})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{g}(\mathbf{y})}\right].$
Let us denote by $\nabla_{\mathbf{g}}h^{(j)}_{\mathbf{y}}(g)$ the row vector of partial derivatives of $\mathbf{h}^{(j)}_{\mathbf{y}}$. Since $\Theta_{d}^{k}$ is finite, then the partial derivatives are bounded, uniformly with respect to $\mathbf{g}\in\Delta^{D}$. Thus, $\mathbf{h}^{(j)}_{\mathbf{y}}(\mathbf{g})$ is uniformly differentiable. By Theorem F.2, $b_{n}^{1/2}(\mathbf{g}_{n}-\tilde{\mathbf{g}})$ converges in the sense of almost sure conditional convergence to a centered normal kernel $\mathcal{N}(0,\mathbf{V}_{\tilde{g}})$, with random covariance matrix $\mathbf{V}_{\tilde{g}}$ as in (F.26). Denote by $D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}})$ the Jacobian matrix of $\mathbf{h}_{\mathbf{y}}$. By Theorem B.5, $b_{n}^{1/2}(\mathbf{h}_{\mathbf{y}}(\mathbf{g}_{n})-\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))$ converges in the sense of almost sure conditional convergence, to a Gaussian kernel with covariance matrix $(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))\mathbf{V}_{\tilde{\mathbf{g}}}(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\tilde{\mathbf{g}}})^{T}$. Its $(j,j^{\prime})$th element can be written as
$((D_{\tilde{\mathbf{g}}}\mathbf{h}_{\mathbf{y}}(\tilde{\mathbf{g}}))\mathbf{V}_{\tilde{\mathbf{g}}}(D_{\tilde{\mathbf{g}}}\mathbf{h}_{\tilde{\mathbf{g}}})^{T})_{j,j^{\prime}}=\hat{\theta}_{\tilde{\mathbf{g}},j}(\mathbf{y})\hat{\theta}_{\tilde{\mathbf{g}},j^{\prime}}(\mathbf{y})$
$\sum_{i,i^{\prime}=1}^{D-1}\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}+\mathbf{e}_{j}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y})}\right]\left[\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})\tilde{g}({\bm{\vartheta}}_{i^{\prime}}\mid\mathbf{z})p_{\tilde{g}}(\mathbf{z})-\tilde{g}({\bm{\vartheta}}_{i})\tilde{g}({\bm{\vartheta}}_{i^{\prime}})\right]$
$\left[\frac{k_{\mathbf{y}+\mathbf{e}_{j^{\prime}}}({\bm{\vartheta}}_{i^{\prime}})-k_{\mathbf{y}+\mathbf{e}_{j^{\prime}}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y}+\mathbf{e}_{j^{\prime}})}-\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})-k_{\mathbf{y}}({\bm{\vartheta}}_{D})}{p_{\tilde{g}}(\mathbf{y})}\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}(\theta_{D})})\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}(\theta_{D})})\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})})\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{i})}-\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})}{\tilde{g}({\bm{\vartheta}}_{D})}-\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}+\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})})\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})})-(1-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{z}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{d}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})(\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})})-(1-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{z}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{d}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}-(1-\tilde{g}({\bm{\vartheta}}_{D}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$\left[\sum_{i=1}^{D-1}\tilde{g}({\bm{\vartheta}}_{i})\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{i})}-(1-\tilde{g}({\bm{\vartheta}}_{D}))\frac{\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{D}\mid\mathbf{y})}{\tilde{g}({\bm{\vartheta}}_{D})}\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]$
$-\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\left[\sum_{i=1}^{D}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\right]\left[\sum_{i=1}^{D}\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y})\right]$
$=\hat{\theta}_{\tilde{g},j}(\mathbf{y})\hat{\theta}_{\tilde{g},j^{\prime}}(\mathbf{y})\sum_{\mathbf{z}\in\mathbb{N}_{0}^{k}}p_{\tilde{g}}(\mathbf{z})\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]\left[\sum_{i=1}^{D}\frac{\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{z})}{\tilde{g}({\bm{\vartheta}}_{i})}(\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}+\mathbf{e}_{j^{\prime}})-\tilde{g}({\bm{\vartheta}}_{i}\mid\mathbf{y}))\right]$
F.3 Proof of Theorem E.2
The proof of Theorem E.2 is based on the following lemmas.
Lemma F.3.
Let $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ be the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{D}$, respectively. Under the assumptions of Theorem E.2, there exists a constant $C=C(d,k,\vartheta_{1},\vartheta_{d})$ such that for every $n\geq 1$
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}.$
Proof.
We can write that
${\rm Regret}(g_{n},g^{*})$
$=\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}(\hat{\theta}_{g_{n},l}(\mathbf{y})-\hat{\theta}_{g^{*},l}(\mathbf{y}))^{2}p_{g^{*}}(\mathbf{y})$
$=\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\left(\frac{\sum_{i=1}^{D}{{\bm{\vartheta}}_{i,l}}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{y_{j}}g_{n}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})}-\frac{\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{y_{j}}g^{*}(\bm{\theta})}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}\right)^{2}$
×$\quad\quad\times\frac{1}{\prod_{j=1}^{k}y_{j}!}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})$
$\leq 2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\left(\sum_{i=1}^{D}e^{-2{\bm{\vartheta}}_{i,l}}{\bm{\vartheta}}_{i,l}^{2y_{l}+1}\prod_{j\neq l}e^{-2{\bm{\vartheta}}_{i,j}}{\bm{\vartheta}}_{i,j}^{2y_{j}}\left(g_{n}({\bm{\vartheta}}_{i})-g^{*}({\bm{\vartheta}}_{i})\right)\right)^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$+2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\left(\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}^{y_{l}+1}e^{-{\bm{\vartheta}}_{i,l}}\prod_{j\neq l}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})\right)^{2}}{\left(\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g_{n}({\bm{\vartheta}}_{i})\right)^{2}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
×$\quad\quad\times\left(\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}(g_{n}({\bm{\vartheta}}_{i})-g^{*}({\bm{\vartheta}}_{i})\right)^{2}$
$\leq 2\sum_{l=1}^{k}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=!}^{k}y_{j}!}\frac{\sum_{i=1}^{D}{\bm{\vartheta}}_{i,l}^{2y_{l}+2}e^{-2{\bm{\vartheta}}_{i,l}}\prod_{j\neq l}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\quad\quad+2\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=!}^{k}y_{j}!}\frac{{\bm{\vartheta}}_{id}^{2}\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\leq 4\vartheta_{d}^{2}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{\prod_{j=1}^{k}y_{j}!}\frac{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{2y_{j}}e^{-2{\bm{\vartheta}}_{i,j}}}{\sum_{i=1}^{D}\prod_{j=1}^{k}{\bm{\vartheta}}_{i,j}^{y_{j}}e^{-{\bm{\vartheta}}_{i,j}}g^{*}({\bm{\vartheta}}_{i})}$
$\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
with $C=4D\vartheta_{d}^{2}exp(k(\vartheta_{d}^{2}/\vartheta_{1}-2\vartheta_{1}+\vartheta_{d})).$
Lemma F.4.
Under the assumptions of Theorem E.2, the stochastic approximation in $\Delta^{D}$
$\mathbf{g}_{n+1}=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$ (F.30)
with $\alpha_{n}>0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\sum_{n=1}^{\infty}\alpha_{n}^{2}<\infty$,
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{k_{\mathbf{Y}_{n+1}}}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\hat{\mathbf{g}}_{n}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right),$
and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}_{G^{*}}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
Proof.
Let $J=\partial\Delta^{D}\cup\{\mathbf{g}^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005) if we can prove that
  • S1
    $\alpha_{n}>0$, $\alpha_{n}\rightarrow 0$, as $n\rightarrow\infty$ and $\sum_{n=1}^{\infty}\alpha_{n}=\infty$,
  • S2
    $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{D}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$,
  • S3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
  • S4
    $\mathbf{h}$ is measurable and locally bounded.
Let $\mathscr{l}$ be defined as
$\mathscr{l}(\mathbf{g})=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\log\left(\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}\right)p_{\mathbf{g}^{*}}(\mathbf{y}).$
S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and S4 is obvious. We now prove S2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{D})}$,
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right)\right)^{T}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}-k_{\mathbf{y},D}\mathbf{1}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)$
$=-\left(\sum_{i=1}^{D}\mathbf{g}_{i}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1\right)^{2}\right)$
$\leq-\left(\sum_{i=1}^{D}\mathbf{g}_{i}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1\right)\right)^{2}=0,$
where we have used Jensen inequality. We now prove that the inequality is strict, unless $\mathbf{g}=\mathbf{g}^{*}$. By contradiction, assume that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ in $\mathring{(\Delta^{D})}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that, when $\mathbf{g}=\mathbf{g}^{\prime}$ and when $\mathbf{g}=\mathbf{g}^{\prime\prime}$, $\sum_{i=1}^{D}\mathbf{g}_{i}(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}-1)^{2}=0$, thus $\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}}(\mathbf{y})}=1$ for every $i=1,\dots,D$. Then, for every $i=1,\dots,D$ for every $a\in(0,1)$
$1=\sum_{i=1}^{D}(a\mathbf{g}^{\prime}({\bm{\vartheta}}_{i})+(1-a)\mathbf{g}^{\prime\prime}({\bm{\vartheta}}_{i}))\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{ap_{\mathbf{g}^{\prime}}(\mathbf{y})+(1-a)p_{\mathbf{g}^{\prime\prime}}(\mathbf{y})}$
$<\sum_{i=1}^{D}(a\mathbf{g}^{\prime}({\bm{\vartheta}}_{i})+(1-a)\mathbf{g}^{\prime\prime}({\bm{\vartheta}}_{i}))\left(a\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}^{\prime}}(\mathbf{y})}+(1-a)\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\frac{p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}^{\prime\prime}}(\mathbf{y})}\right)$
$=1,$
which leads to a contradiction. Hence S2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g})=\sum_{i=1}^{D}\log(\mathbf{g}^{*}_{i}/\mathbf{g}_{i})\mathbf{g}^{*}_{i}$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$. Let $\tilde{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})+\nabla_{\mathbf{g}}\mathcal{L}(\mathbf{g}_{n})^{T}(\mathbf{g}_{n+1}-\mathbf{g}_{n})$
$+(\mathbf{g}_{n+1}-\mathbf{g}_{n})^{T}\nabla^{2}_{\mathbf{g}}\mathcal{L}(\tilde{\mathbf{g}}_{n})(\mathbf{g}_{n+1}-\mathbf{g}_{n}).$
Then,
${\mathcal{L}}(\mathbf{g}_{n+1})$ $=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Notice that
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
with
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$
and
$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E_{G^{*}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
$>\liminf_{n}E_{G^{*}}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}_{G^{*}}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E_{G^{*}}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $\mathbf{Y}_{1},\dots,\mathbf{Y}_{n}$, we can write that
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}}=1}^{D}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{(\sum_{{i^{\prime\prime}}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})\mathbf{g}_{n}({\bm{\vartheta}}_{i^{\prime\prime}}))^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}^{*}({\bm{\vartheta}}_{i^{\prime}})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\sum_{i,{i^{\prime}},{i^{\prime\prime}}=1}^{D}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}^{*}({\bm{\vartheta}}_{i^{\prime}})\mathbf{g}_{n}({\bm{\vartheta}}_{i^{\prime\prime}})\right)$
$\leq 2\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}\left(1+D\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\max_{i,{i^{\prime}},{i^{\prime\prime}}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})\right)$
$=o(\alpha_{n}),$
where the last equality holds since for every $i,{i^{\prime}},{i^{\prime\prime}}$
$\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})^{2}}{k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime}})^{2}}k_{\mathbf{y}}({\bm{\vartheta}}_{i^{\prime\prime}})$ $=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{\prod_{j=1}^{k}e^{-2{\bm{\vartheta}}_{i,j}}\vartheta_{i,j}^{2\mathbf{y}_{j}}/(\mathbf{y}_{j}!)^{2}}{\prod_{j=1}^{k}e^{-2{\bm{\vartheta}}_{i^{\prime},j}}\vartheta_{i^{\prime},j}^{2\mathbf{y}_{j}}/(\mathbf{y}_{j}!)^{2}}\prod_{j=1}^{k}e^{-{\bm{\vartheta}}_{i^{\prime\prime},j}}\vartheta_{i^{\prime\prime},j}^{\mathbf{y}_{j}}/\mathbf{y}_{j}!$
$=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\prod_{j=1}^{k}e^{2{\bm{\vartheta}}_{i^{\prime},j}-2{\bm{\vartheta}}_{i,j}-{\bm{\vartheta}}_{i^{\prime\prime},j}}\left(\frac{{\bm{\vartheta}}_{i,j}^{2}{\bm{\vartheta}}_{i^{\prime\prime},j}}{{\bm{\vartheta}}_{i^{\prime},j}^{2}}\right)^{y}\frac{1}{y_{j}!}$
$=\prod_{j=1}^{k}\sum_{\mathbf{y}_{j}\in\mathbb{N}_{0}}e^{2{\bm{\vartheta}}_{i^{\prime},j}-2{\bm{\vartheta}}_{i,j}-{\bm{\vartheta}}_{i^{\prime\prime},j}}\left(\frac{{\bm{\vartheta}}_{i,j}^{2}{\bm{\vartheta}}_{i^{\prime\prime},j}}{{\bm{\vartheta}}_{i^{\prime},j}^{2}}\right)^{y}\frac{1}{y_{j}!}<\infty.$
It follows that,
$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
Lemma F.5.
Under the assumptions of Theorem 2.4, for every $\delta<2-1/\gamma$,
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
Proof.
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P_{G^{*}}$-a.s. to $\mathbf{g}\in\Delta^{D}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{D}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{D}$:
$\mathbf{g}_{n+1}$ $=\mathbf{g}_{n}+\alpha_{n+1}\mathbf{h}(\mathbf{g}_{n})+\alpha_{n+1}{\bm{\epsilon}}_{n+1},$
with initial value $\mathbf{g}_{0}$, where
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})-\bf 1\right),$
and
${\bm{\epsilon}}_{n+1}=\mathbf{g}\circ\left(\frac{k_{\mathbf{Y}_{n+1}}}{p_{\mathbf{g}}(\mathbf{Y}_{n+1})}p_{\mathbf{g}^{*}}(\mathbf{Y}_{n+1})-\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}}{p_{\mathbf{g}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right).$
The thesis follows from Theorem 3.1.1 in Chen (2005) if we can prove that
  • T1
    $\alpha_{n}\rightarrow 0$, $\sum_{n=1}^{\infty}\alpha_{n}=\infty$, $\alpha_{n+1}^{-1}-\alpha_{n}^{-1}\rightarrow 0$,
  • T2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.
  • T3
    $\mathbf{h}$ is measurable and locally bounded, and is differentiable at $\mathbf{g}^{*}$. Let $H$ be a matrix such that, as $\mathbf{g}\rightarrow\mathbf{g}^{*}$,
    $\mathbf{h}(\mathbf{g})=\mathbf{H}(\mathbf{g}-\mathbf{g}^{*})+\mathbf{r}(\mathbf{g}),\quad\mathbf{r}(g^{*})=0,\quad\mathbf{r}(g)=o(||\mathbf{g}-\mathbf{g}^{*}||).$
    All the eigenvalues of the matrix $H$ have negative real parts.
Remark F.6.
Theorem 3.1.1 in Chen (2005) also requests the following condition:
T4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma F.4.
B1 is obvious. To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale. Moreover, there exoists a constant $C$ such that, for every $i=1,\dots,D$,
$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i}^{2})$
$\leq C\sum_{k=1}^{\infty}\left(\frac{\alpha}{\alpha+k}\right)^{\gamma(2-\delta)}<\infty,$
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and
$\mathbf{H}_{i,j}=-\mathbf{g}^{*}_{i}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}.$
This matrix can be seen as the product of the diagonal matrix ${\rm diag}(\mathbf{g}^{*})$, which is positive definite, times $-\mathbf{M}$, with $\mathbf{M}_{i,j}=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}$. We prove that $M$ is positive definite. Given a vector $\mathbf{u}\in\mathbb{R}^{D}$ with $||\mathbf{u}||=1$, we can write that
$\mathbf{u}^{T}\mathbf{M}\mathbf{u}=$ $\sum_{i,j=1}^{D}\mathbf{u}_{i}\mathbf{u}_{j}\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})k_{\mathbf{y}}({\bm{\vartheta}}_{j})}{p_{\mathbf{g}^{*}}(\mathbf{y})}=\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{1}{p_{\mathbf{g}^{*}}(\mathbf{y})}\left(\sum_{i=1}^{D}\mathbf{u}_{i}k_{\mathbf{y}}({\bm{\vartheta}}_{i})\right)^{2}\geq 0,$
with equality if and only if $\sum_{i=1}^{D}\mathbf{u}_{i}k_{\mathbf{y}}({\bm{\vartheta}}_{i})=0$ for every $\mathbf{y}\in\mathbb{N}_{0}^{k}$. The only solution to the last equation in $\mathbf{u}$ is the zero vector, which is inconsistent with $||\mathbf{u}||=1$. Since the function $\mathbf{u}^{T}\mathbf{M}\mathbf{u}$ is continuous on the compact set $||\mathbf{u}||=1$, then the minimum is attained, and by the above reasoning it is strictly positive. It follows that $\mathbf{M}$ is positive definite. ∎
of Theorem E.2.
Denoting by $(\mathbf{g}_{n})$ and $\mathbf{g}^{*}$ the representations of $(g_{n})$ and $g^{*}$ in $\mathbb{R}^{D}$, we can write, by Lemma F.3 and Lemma F.5,
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta}).$